### v1

# The number of cultural traits, evolving genealogies, and the descendant process

Abstract
We consider a Moran-type model of cultural evolution, which describes how traits emerge, are transmitted, and get lost in populations. Our analysis focuses on the underlying cultural genealogies; they were first described by Aguilar and Ghirlanda (2015) and are closely related to the ancestral selection graph of population genetics, wherefore we call them ancestral learning graphs. We investigate their dynamical behaviour, that is, we are concerned with evolving genealogies. In particular, we consider the total length of the genealogy of a sample of individuals from a stationary population as a function of the (forward) time at which the sample is taken. This quantity shows a sawtooth-like dynamics with linear increase interrupted by collapses to near-zero at random times. We relate this to the metastable behaviour of the stochastic logistic model, which describes the evolution of the number of ancestors, or equivalently, the number of descendants of a given sample.
We assume that new inventions appear independently in every individual, and all traits of the cultural parent are transmitted to the learner in any given learning event. The set of traits of an individual then agrees with the set of innovations along its genealogy. The properties of the genealogy thus translate into the properties of the trait set of a sample. In particular, the moments of the number of traits are obtained from the moments of the total length of the genealogy.
keywords:
cultural evolution , ancestral learning graph , ancestral selection graph , stochastic logistic model , metastability , coalescence
journal: Theoretical Population Biology
[…]
[1]organization=School of Interdisciplinary Mathematical Sciences, addressline=4-21-1 Nakano, Nakano, postcode=164-8525, city=Tokyo, country=Japan
[…]
[2]organization=Research Center for Integrative Evolutionary Science, SOKENDAI (The Graduate University for Advanced Studies), addressline=Shonan Village, Hayama, postcode=240-0193, city=Kanagawa, country=Japan
[…]
[3]organization=Kochi University of Technology, addressline=2-22 Eikokuji, Kochi City, postcode=780-8515, city=Kochi, country=Japan
[…]
[4]organization=Faculty of Technology, addressline=Bielefeld University, postcode=33501, city=Bielefeld, country=Germany
1 Introduction
Stochastic processes in cultural evolution describe how traits emerge, are transmitted, and get lost in populations under certain random laws. This relatively new topic of research bears many connections with population genetics with respect to the models used, the questions asked, and the methods employed. For example, Strimling et al. (2009), Fogarty et al. (2015, 2017), and Aoki (2018) studied a discrete-time Moran model of cultural evolution under various assumptions on the origination and transmission of traits; they investigated quantities such as the number of traits maintained in a population and the corresponding popularity spectrum, that is, the histogram of trait frequencies in the population, akin to the site-frequency spectrum in population genetics. These and similar results are centered around the traits, focus on the stationary state, and characterise it by means of expectations of various quantities related to trait frequencies.
Aguilar and Ghirlanda (2015) were, to the best of our knowledge, the first to explicitly consider the cultural genealogies in this kind of model. They worked with a continuous-time Moran-type model of cultural evolution, where the genealogy is closely related to the ancestral selection graph (ASG) of population genetics (Neuhauser and Krone, 1997; Krone and Neuhauser, 1997). Aguilar and Ghirlanda (2015) investigated the expected time to the most recent unique ancestor (MRUA), a quantity different from both the most recent common ancestor (MRCA) and the ultimate ancestor (UA) of the ASG. Kobayashi et al. (2018) worked with a discrete-time Wright–Fisher-type model of cultural evolution assuming independent origination and transmission of traits and used genealogical thinking to obtain the expected number and age of distinct cultural traits in a finite sample. Altogether, little is known beyond expectations.
The goal of this paper is threefold. First, we will focus on the underlying cultural genealogies, which are also of independent interest. Second, we aim at more details about the (stationary) distribution of the number of traits; that is, we look at the variance and higher moments. Third, we consider the temporal fluctuations of the number of traits at stationarity. All properties of the trait frequencies are closely tied to the genealogies as well. Exploiting the connection to recent variants of the ASG of population genetics and to the stochastic logistic model of epidemiology and theoretical ecology, we analyse both the mean behaviour and the dynamics. A special role will be played by the concept of evolving genealogies and the complementary descendant process.
The paper is organised as follows. In Section 2, we describe and define our model forward in time and then introduce the genealogical process, which we term the ancestral learning graph (ALG), and elaborate on the connection to ancestral structures in population genetics. In Section 3, we present the results. We first show that the moments of the number of traits in a sample at stationarity can be expressed in terms of the moments of the total tree length of the ALG, that is, the sum of the lengths of all its branches. For the mean and variance, we derive explicit expressions. In simulations, we explore the dynamics of the tree length (and hence the number of traits) as a function of time. We observe an unexpeced sawtooth-like behaviour, which cannot be explained by the moments alone. We understand it by formulating the concept of evolving genealogies (backward in time) and the descendant process (forward in time) and investigating them with the help of the properties of the stochastic logistic process. In particular, the metastable state of the latter and the coalescence of sets of lines in the ALG will play crucial roles. As an aside, we also obtain the popularity spectrum. We discuss our findings in Section 4.
2 The model
We follow the model of Aguilar and Ghirlanda (2015), who assume a population of fixed size $N$ in continuous time with death-birth and learning events, akin to a Moran model in population genetics. Newborn individuals do not have any cultural parent at the time of birth, but acquire role models one by one through learning events, each of which assigns a (single) new cultural parent to the focal individual; it is intentionally left unspecified which traits are transmitted in a given learning event, and whether and how new traits emerge. We combine this model with a simplistic model that makes the appearance and transmission of traits explicit. This is motivated by the work of Kobayashi et al. (2018), who assume a discrete-time Wright–Fisher process, where every newborn invents a random number of new traits and is assigned a random number $K$ of potential role models; every trait of each of the role models is independently transmitted to the newborn with probability $b$. Here we assume that, within the continuous-time model of Aguilar and Ghirlanda (2015), new traits appear at constant rate $\mu$ in every individual, and we specify the mode of transmission in accordance with Kobayashi et al. (2018) with the choice $b=1$, so that the learning individual acquires all traits of its cultural parent(s). The resulting model may also be considered as a continuous-time version of (special cases of) the models of Strimling et al. (2009), Fogarty et al. (2015, 2017), and Aoki (2018).
2.1 The forward process
Consider a population of $N$ individuals. From the Moran model, we borrow the graphical representation shown in Figures 1 and 2. Each individual corresponds to a vertical line segment, with the forward direction of time being top to bottom; the lines have labels in $[N]\mathrel{\mathop{:}}=\{1,2,\ldots,N\}$. The events described below are represented by graphical elements juxtaposed to this picture. Again in line with a common strategy for the Moran model, we describe the model in two steps: first the so-called untyped model, which contains the death-birth and learning events; and second the typed model, which includes the traits, specifically, the innovation events and the transmission of traits on the occasion of learning events.
  1. 1.
    Untyped version. Each individual can experience two kinds of events at any instant of time and independently of the remaining population: a death event and a learning event. Death occurs at rate $u>0$ per individual, and when this happens, a newborn will replace it. Learning events occur at rate $s>0$ per individual. In a learning event, the individual randomly samples a cultural parent uniformly from among the $N$ individuals (including itself). So far, the learning events are abstract; we do not yet say what is transmitted along them. In the graphical representation (Fig. 1), death-birth events are symbolised by crosses; they appear at rate $u$ on every line, by way of independent Poisson point processes. Learning events are depicted by arrows between the lines, with the cultural parent at the tail and the learner at the tip; arrows appear at rate $s/N$ per ordered pair of lines, again by way of independent Poisson point processes. We omit arrows where the parent and the learner are identical.
    Figure 1: Graphical representation of a realisation of the untyped model ($N=5$).
  2. 2.
    Including the types. The type of an individual is the collection of its cultural traits. Given a realisation of the untyped system, we turn it into a typed one as folllows. We assume that every individual independently invents new traits at rate $\mu>0$. In the graphical representation, these innovations are indicated by circles, which are laid down on every line by way of a Poisson point process at rate $\mu$, see Figure 2. We assume that, first, all innovations are non-recurrent, that is, have never occurred in the population before. Second, in every learning event, the learner acquires all traits carried by the cultural parent. Third, we assume that death events eliminate all traits, so the newborn is devoid of any trait.
    We now assign a set of traits to each line at time $t=0$ in an exchangeable way (that is, according to a law that is invariant under permutation of lines); this set may or may not be empty. If we then propagate the traits forward in time according to the rules just described, we get the types, that is, the collection of traits, for every individual and any time $t>0$.
Figure 2: Graphical representation of a realisation of the typed model.
2.2 Formal definition of the forward process
Let
$\Phi(t)\mathrel{\mathop{:}}=(K_{1}(t),\ldots,K_{N}(t),\Omega(t))$ (1)
be the state of the typed graphical construction at time $t$, where $K_{\alpha}(t)\subset\mathbb{N}$ denotes the set of traits that individual $\alpha\in[N]$ knows at time $t\in\mathbb{R}_{\geq 0}$, and $\Omega(t)\in\mathbb{N}_{0}$ denotes the number of traits ever invented until $t$, including those already present at time 0; so $\Omega(0)=\lvert\cup_{\alpha\in[N]}K_{\alpha}(0)\rvert$ and $K_{\alpha}(t)\subseteq\{1,2,\dots,\Omega(t)\}$ (where the latter is agreed to be $\varnothing$ if $\Omega(t)=0$), and
$\underbrace{2^{\mathbb{N}}\times\dots\times 2^{\mathbb{N}}}_{N\text{ times}}\times\mathbb{N}_{0}$ (2)
is the state space. Note that $\Omega(t)=\lvert\cup_{t^{\prime}\in[0,t]}\cup_{\alpha\in[N]}K_{\alpha}(t^{\prime})\rvert$. The process starts at $\Phi(0)=(K_{1}(0),\ldots,K_{N}(0),\Omega(0))$ and evolves as follows. If $\Phi(t)=(k_{1},\ldots,k_{N},\omega)$, the following events may happen:
At rate $Nu$, a death event occurs and a uniformly-chosen individual $\alpha\in[N]$ loses all traits, so $k_{\alpha}$ changes to $\varnothing$, and we see the transition
$(k_{1},\ldots,k_{\alpha},\ldots,k_{N},\omega)\longrightarrow(k_{1},\ldots,\varnothing,\ldots,k_{N},\omega)$ (3a)
with $\varnothing$ in position $\alpha$. At rate $N\mu$, an innovation event occurs and a uniformly-chosen individual $\alpha$ acquires a new trait, so $\omega$ changes to $\omega+1$, $k_{\alpha}$ changes to $k_{\alpha}\cup\{\omega+1\}$, and the transition is
$(k_{1},\ldots,k_{\alpha},\ldots,k_{N},\omega)\longrightarrow(k_{1},\ldots,k_{\alpha}\cup\{\omega+1\},\ldots,k_{N},\omega+1).$ (3b)
At rate $Ns$, a learning event occurs and a uniformly-chosen individual $\alpha$ learns from another uniformly-chosen individual $\beta$, so $k_{\alpha}$ is replaced by $k_{\alpha}\cup k_{\beta}$ and
$(k_{1},\ldots,k_{\alpha},\ldots,k_{N},\omega)\longrightarrow(k_{1},\ldots,k_{\alpha}\cup k_{\beta},\ldots,k_{N},\omega);$ (3c)
note that nothing happens if $\alpha=\beta$. Note also that we have defined the process in a way that is close to how we will later simulate it; in this formulation, it does not become stationary for $t\to\infty$ since, due to the consecutive labelling of the traits, both $(\Omega(t))_{t\geq 0}$ and the $(K_{\alpha}(t))_{t\geq 0}$ are transient. The quantities we will analyse, however, will only rely on the number of traits rather than their labels and do become stationary, as will become obvious in the backward picture.
For a subset of individuals $G\subseteq[N]$, we define their knowledge set as
$K_{G}(t):=\bigcup_{\alpha\in G}K_{\alpha}(t)$ (4)
by slight abuse of notation (so $K_{\{\alpha\}}(t)=K_{\alpha}(t)$). With this, we define the number of traits of a ‘typical’ group of $n$ individuals at time $t$ as
$C_{n}(t)=|K_{[n]}(t)|.$ (5)
By ‘typical’, we here allude to the fact that, by exchangeability, the $K_{G}(t)$ are identically distributed for all $G\subseteq[N]$ with $\lvert G\rvert=n$, so we choose $G=[n]$ as their representative. We will see below that the distribution of $(C_{n}(t))_{t\geq 0}$ will become stationary as $t\to\infty$; we will denote by $C_{n}=C_{n}(\infty)$ a random variable that has this stationary distribution.
2.3 The ancestral learning graph (ALG)
Let us now take a backward perspective and consider the genealogy of $n\in[N]$ individuals sampled at forward time $t$, to which we will refer as the present; we may, but need not, choose $t$ as some fixed final time $t_{\text{max}}$. Starting again from the untyped version of the graphical representation, let us trace back the ancestral lines of the sample, as illustrated in Figure 3. More precisely, we first describe the untyped version of the ancestral learning graph (ALG), which consists, at any given time, of the set of all cultural parents, parents of parents and so forth, that is, of all cultural ancestors, of the sample. It is constructed as follows.
Figure 3: The ALG for the realisation of the forward model in Figure 2, starting from $t=t_{\text{max}}$. Red and black lines are ancestral and non-ancestral, respectively, to the sample of size $n=2$ taken at $\tau=0$.
  1. 1.
    Start the graph by tracing back the lines emerging from the individuals sampled at forward time $t$. Denote backward time by $\tau$, so backward time $\tau$ corresponds to forward time $t-\tau$; in particular, forward times 00 and $t$ correspond to backward times $\tau=t$ and $\tau=0$, respectively. Proceed as follows in an iterative way in the backward direction of time until the initial time is reached, or until there are no lines left.
    1. (a)
      If a line currently in the graph is hit by the tip of a learning arrow, we trace back both ancestors, namely that of the cultural parent (at the tail of the arrow, also known as the incoming branch) and that of the learning individual (at the tip; also known as the continuing branch). That is, we add the incoming line to the graph; this results in a branching event. Note that the graph remains unchanged if the incoming branch is already contained in it.
    2. (b)
      If a line currently in the graph is hit by a cross, we do not follow it back any further, that is, we prune it.
    The resulting untyped ALG consists of all cultural ancestors of the sampled individuals over time, including the individuals in the sample themselves until they die (that is, individuals are counted as their own cultural ancestors); the untyped ALG corresponds to the untyped version of the forward model.
  2. 2.
    The untyped ALG may be turned into a typed one by first assigning, to each line in the graph that is still alive at backward time $t$, the initial trait set from the Moran model; if there are no lines left at backward time $t$, this step is void. Then, the traits are propagated forward (that is, downward) in the genealogy according to the same rules as before. That is, when a new trait was acquired in the Moran model on a line belonging to the genealogy, it is added to the trait set of the individual; in each branching event, the trait set of the parent is added to that of the learner; if a line encounters a cross, it loses all traits. This way, a trait set is associated with every line element of the graph.
Note that the untyped ALG obtained in Step 1 is the cultural genealogy of Aguilar and Ghirlanda (2015), with the only difference that we allow for individuals to also choose themselves as cultural parents; but this is an event without any effect, so it only introduces a tiny change of time scale. The typing that happens in Step 2 describes how traits appear and are transmitted along the given genealogy; this step was intentionally left unspecified by Aguilar and Ghirlanda (2015).
2.4 The ALG as a stochastic process
So far, we have described the ALG that results when a realisation of the forward model is given. We now define it as a stochastic process independent of such a realisation, again first in the untyped version denoted by $(\Lambda(\tau))_{\tau\geq 0}$, and then in the typed version.
  1. 1.
    For the untyped version, let $\Lambda_{\alpha}(\tau)\subseteq[N]$ denote the set of ancestors at backward time $\tau$ of individual $\alpha\in[N]$ at backward time 00; and, for $G\subseteq[N]$, set $\Lambda_{G}(\tau)\mathrel{\mathop{:}}=\cup_{\alpha\in G}\Lambda_{\alpha}(\tau)$ by slight abuse of notation (so $\Lambda_{\{\alpha\}}(\tau)=\Lambda_{\alpha}(\tau)$). Then the process $(\Lambda(\tau))_{\tau\geq 0}\mathrel{\mathop{:}}=\big{(}\Lambda_{1}(\tau),\ldots,\Lambda_{N}(\tau)\big{)}_{\tau\geq 0}$ contains the complete genealogical information of all individuals in the present population. It evolves as follows. Clearly $(\Lambda_{1}(0),\ldots,\Lambda_{N}(0))=(\{1\},\{2\},\ldots,\{N\})$. If, for any time $\tau\geq 0$, the current state is $\Lambda(\tau)=(\lambda_{1},\ldots,\lambda_{N})$, the following events may happen: at rate $Nu$, a death event occurs and a randomly-chosen individual $\beta$ dies. So $\beta$ is removed from the set of ancestors of anybody, and
    $(\lambda_{1},\ldots,\lambda_{N})\longrightarrow(\lambda_{1}\setminus\{\beta\},\ldots,\lambda_{N}\setminus\{\beta\}).$ (6a)
    At rate $Ns$, a learning event occurs and a randomly-chosen individual $\beta_{1}$ learns from another randomly chosen individual $\beta_{2}$ (where $\beta_{2}=\beta_{1}$ is allowed, in which case the event is silent). So if, for $\alpha\in[N]$, $\beta_{1}$ is contained in $\lambda_{\alpha}$, then $\beta_{2}$ joins this ancestral set, and otherwise nothing happens:
    $\lambda_{\alpha}\longrightarrow\begin{cases}\lambda_{\alpha}\cup\{\beta_{2}\},&\text{if }\beta_{1}\in\lambda_{\alpha},\\ \lambda_{\alpha},&\text{otherwise}\end{cases}\qquad\alpha\in[N].$ (6b)
    In words, any individual $\alpha$ who has $\beta_{1}$ as its ancestor just before the focal learning event will add $\beta_{2}$ as another ancestor as a result of the event.
  2. 2.
    The untyped ALG may be turned into a typed one by superposing the untyped graph with a Poisson point process that lays down innovations at rate $\mu$ on every line. Together with an initial assignment of types to the lines alive at backward time $t$ and propagation along the graph as in the forward model, this determines the types along all lines in the graph.
Returning to the untyped version, let us note for later use that, since $\Lambda_{\alpha}(\tau)\subseteq\Lambda_{[N]}(\tau)$,
[…]
for any $\alpha\in[N]$.
In what follows, an important role will be played by the line-counting process $(Y_{n}(\tau))_{\tau\geq 0}$ of the ALG. That is,
$Y_{n}(\tau)\mathrel{\mathop{:}}=\lvert\Lambda_{[n]}(\tau)\rvert$ (8)
is the number of lines at backward time $\tau$ of the sample $[n]$ of individuals taken at backward time 00; due to exchangeability, we allow ourselves to simply speak of a sample of size $n$. Also, we will sometimes omit the dependence on the initial value $n$. $(Y(\tau))_{\tau\geq 0}$ is a birth-death process [^1] This birth-death process is not to be confused with the death-birth events in the model, as indicated by crosses in the graphical representation. on $[N]_{0}:=\{0,1,\ldots,N\}$ with birth and death rates222not to be confused with the realisations $\lambda_{\alpha}$ of $\Lambda_{\alpha}$ and the innovation rate $\mu$.
$\lambda_{y}\mathrel{\mathop{:}}=sy\frac{N-y}{N}\quad\text{and}\quad\mu_{y}\mathrel{\mathop{:}}=uy,$ (9)
respectively, when $Y(\tau)=y$ (note that the factor $(N-y)/N$ is the probability that a learning arrow comes from outside the current graph). Clearly, $(Y(\tau))_{\tau\geq 0}$ is an absorbing Markov chain with 0 as the only absorbing state. The behaviour of the process is very well studied, since it is, at the same time, the stochastic logistic process of ecology and epidemiology; in particular, $(Y(\tau))_{\tau\geq 0}$ is the number of infected individuals in a stochastic SIS model with infection rate $s$ and recovery rate $u$ per individual (Andersson and Britton, 2000, Ch. 8.2).
Another crucial quantity will be $L_{n}$, the tree length, that is, the total length of all branches, as $\tau\to\infty$, in the ALG of a sample of size $n$, that is,
$L_{n}\mathrel{\mathop{:}}=\lim_{\tau\to\infty}L_{n}(\tau),\quad\text{where}\quad L_{n}(\tau)\mathrel{\mathop{:}}=\int_{0}^{\tau}Y_{n}(r)\mathrm{d}{r}=\int_{0}^{\tau}\lvert\Lambda_{[n]}(r)\rvert\mathrm{d}r.$ (10)
Indeed, the limit exists and is finite for almost all realisations of the ALG, since $(Y_{n}(\tau))_{\tau\geq 0}$ absorbs in 0 almost surely in finite time for any given $n$. In epidemiology, the quantity $L_{n}$, sometimes up to a constant, is referred to as the “cost of an epidemic”, because it measures the total time that individuals spend in the infected state in one episode of the epidemic started with $n$ individuals (Jerwood, 1970; McNeil, 1970; Downton, 1972; Gani and Jerwood, 1972; Ball, 1986; Hernandez-Suarez and Castillo-Chavez, 1999; Andersson and Britton, 2000, Secs. 2.2 and 2.4; Crawford et al., 2018). Previous work has concentrated on its mean $\mathbb{E}[L_{n}]$ or its Laplace-transform $\mathbb{E}[e^{-\theta L_{n}}]$ (or its moment-generating function $\mathbb{E}[e^{\theta L_{n}}]$) but, to the best of our knowledge, explicit expressions for the variance (or higher-order moments) are unavailable so far (but see Stefanov and Wang (2000) for a numerical evaluation of the variance).
2.5 The connection with population genetics
Let us point out the close connections with population genetics theory. The untyped model is closely related with an untyped Moran model with both neutral and selective reproduction events; for review, see Baake and Wakolbinger (2018). There are two main differences: first, cultural traits are not genetically inherited; we therefore do not keep track of the biological parent of a new individual, so the neutral reproduction events of the Moran model turn into death-birth events that only involve a single line. Second, in a learning event, both the incoming and the continuing branches are true cultural parents (recall that learners are counted as their own cultural parents), whereas, in population genetics, they are potential parents, only one of which is the true (genetic) parent.
In the same vein, the untyped ALG is closely related to the ASG of Neuhauser and Krone (1997), which contains all potential ancestors of a sample of individuals. More precisely, in our finite-$N$ setting, the untyped ALG is a special case of the killed ASG of Baake et al. (2023), but here without multiple branching, without coalescence events, and without killing. In the typed model, the innovation mechanism is similar to the infinite-sites model of population genetics (see, for example, Wakeley (2009, Ch. 1.2)), where sequences of infinite length are considered and every mutation hits a site that has never mutated before.
If we reduce the number of possible traits to one and do not allow for innovations (that is, consider the limiting case $\mu=0$), we only have two types of individuals: those that have and those that do not have the trait. Let $X(t)\in[N]_{0}$ be the number of individuals with the trait (so $N-X(t)$ do not have it). If $X(t)=x$, we have a transition to $x-1$ at rate $ux$ (death) and to $x+1$ at rate $sx(N-x)/N$ (learning). This is equivalent to a two-type Moran model with selective reproduction at rate $s$ and deleterious mutation at rate $u$, but without neutral and without frequency-dependent reproduction events; that is, a special case of the model tackled by Baake et al. (2023). Furthermore, note that $(X(t))_{t\geq 0}$ has the same birth and death rates as $(Y(\tau))_{\tau\geq 0}$ and hence the same law.
It has been shown by Baake et al. (2023, Theorem 2.3) that the type-frequency process of the Moran model with mutation and frequency-dependent selection on the one hand and the line-counting process of the killed ASG on the other hand are in factorial moment duality with each other. This translates into the factorial moment duality between the type-frequency process of the above single-trait learning model and the line-counting process of the corresponding untyped ALG as follows. For $z,m\in\mathbb{N}_{0}$, let
$z^{\underline{m}}\mathrel{\mathop{:}}=\begin{cases}1,&m=0,\\ z(z-1)\cdots(z-m+1),&m\geq 1\end{cases}$ (11)
denote the falling factorial. The processes $(X(t))_{t\geq 0}$ and $(Y(\tau))_{\tau\geq 0}$ are dual with respect to the duality function
$H(x,y)\mathrel{\mathop{:}}=\frac{{(N-x)}^{\underline{y}}}{{N}^{\underline{y}}},\quad x,y\in[N]_{0},$ (12)
that is, $(X(t))_{t\geq 0}$ and $(Y(\tau))_{\tau\geq 0}$ satisfy the relation
$\mathbb{E}[H(X(t),y)\mid X(0)=x]=\mathbb{E}[H(x,Y(t))\mid Y(0)=y]$ (13)
for $x,y\in[N]_{0}$ and $t\geq 0$.
Note that $H(x,y)$ is the probability to obtain only individuals without the trait when sampling $y$ individuals without replacement from a population that contains $x$ individuals with the trait (and $N-x$ individuals without it); and the duality means that the sampling probability at time $t$ can be obtained either via the forward or via the backward process. As a consequence, one may obtain, and gain insight into, properties of the learning model forward in time by studying its dual process. See Möhle (1999) for general background on duality in models of population genetics. Since $(X(t))_{t\geq 0}$ and $(Y(\tau))_{\tau\geq 0}$ have the same law, they are actually self dual with respect to $H$.
2.6 Law of large numbers
Let us briefly comment on the deterministic limit, where $N\to\infty$ without rescaling of parameters or time. Then $(Y(\tau))_{\tau\geq 0}$ turns into a linear birth-death process $(\widetilde{Y}(\tau))_{\tau\geq 0}$ on $\mathbb{N}_{0}$ with birth rate $s$ and death rate $u$ per individual. In contrast to the finite-$N$ case, $(\widetilde{Y}(\tau))_{\tau\geq 0}$ does not necessarily absorb in 0. For $s\leq u$, $(\widetilde{Y}(\tau))_{\tau\geq 0}$ is (sub)critical with $\mathbb{P}(\lim_{\tau\to\infty}\widetilde{Y}(\tau)=0\mid\widetilde{Y}(0)=1)=1$; for $s>u$, $(\widetilde{Y}(\tau))_{\tau\geq 0}$ is supercritical, and $\mathbb{P}(\lim_{\tau\to\infty}\widetilde{Y}(\tau)=0\mid\widetilde{Y}(0)=1)=u/s<1$. If the process does not die out, it grows to infinite size almost surely. (These are classical results from the theory of branching processes (Athreya and Ney, 1972, Ch. III.4).)
In the deterministic limit of the single-trait model, the sequence of processes $(X^{(N)}(t)/N)_{t\geq 0}$ (where the upper index indicates the population size) converges, as $N\to\infty$, to the solution of the initial value problem
$\dot{\xi}(t)=\xi(t)[s(1-\xi(t))-u],\quad\xi(0)=\xi_{0}\in[0,1],$ (14)
provided $X^{(N)}(0)/N\to\xi_{0}$. The differential equation has two equilibria: one at $1-u/s$, the other at 00. They perform a transcritical (or exchange of stability) bifurcation at $s=u$: for $s<u$, the equilibrium at 00 is attracting, while the one at $1-u/s$ (which is then $<0$) is repelling; and vice versa for $s>u$. For $s=u$, the two equilibria coincide at 0, which is then attracting.
The duality (13) carries over to the deterministic limit; specifically, evaluating (13) for $y=1$, $N\to\infty$ such that $x/N\to\xi_{0}$, and $t\to\infty$ gives for the unique stable equilibrium $\bar{\xi}$ of the differential equation that $1-\bar{\xi}=\mathbb{E}[(1-\xi_{0})^{\widetilde{Y}(\infty)}\mid\widetilde{Y}(0)=1]$ for $\xi_{0}\in(0,1]$ and so $\bar{\xi}=1-\mathbb{P}(\lim_{\tau\to\infty}\widetilde{Y}(\tau)=0\mid\widetilde{Y}(0)=1)$. For $s\leq u$, therefore, the trait is always absent at equilibrium, while, for $s>u$, the trait is present in a positive proportion of individuals at the stable equilibrium. See Baake et al. (2018) or Baake and Wakolbinger (2018) for the details about the deterministic limit.
Let us note for later use that the differential equation (14) is not only the deterministic limit of our single-trait model, but also the deterministic limit of $(Y^{(N)}(\tau)/N)_{\tau\geq 0}$ (that is, for $N\to\infty$ in the sequence of processes $(Y^{(N)}(\tau)/N)_{\tau\geq 0}$ with population size $N$, without rescaling of parameters or time), provided that $Y^{(N)}(0)/N$ converges; this is clear because $(X(t))_{t\geq 0}$ and $(Y(t))_{t\geq 0}$ have the same law.
In the following, we continue to adhere to finite $N$, but the two regimes $s<u$ and $s>u$ still behave in qualitatively different ways (as long as $s$ is not too close to $u$ and with a smooth transition between the two regimes), see Nåsell (2011, Chap. 2.5) or Foxall (2021). We will therefore continue to use the notions subcritical and supercritical, in line with the literature.
3 Results
One of our major aims is to study the dynamics and the stationary distribution of $(C_{n}(t))_{t\geq 0}$. The typed ALG is the appropriate genealogical structure for this. Since it results from the untyped one by superposition of a Poisson point process that lays down innovations at rate $\mu$ on every line, and every innovation on the tree is passed on to the sample by construction, the crucial quantity is the tree length $L_{n}$ of (10). (Indeed, since we consider $t\to\infty$, the relevant tree length is $\lim_{\tau\to\infty}L_{n}(\tau)$.) By the genealogical picture,
$C_{n}\sim\mathrm{Poi}(\mu L_{n}),$ (15)
that is, $C_{n}$ follows the Poisson distribution with random parameter $\mu L_{n}$. Here and in what follows (and in line with the standard monographs (Durrett, 2008; Wakeley, 2009) in mathematical population genetics and coalescent theory), we let the genealogical picture speak for itself. Our first goal now is to characterise the distribution of $C_{n}$ by calculating moments of $C_{n}$ of any order.
3.1 Calculation of moments
3.1.1 Number of cultural traits
The $m$-th factorial moment of a random variable $W\sim\mathrm{Poi}(\nu)$ is $\mathbb{E}\left[W^{\underline{m}}\right]=\nu^{m}$; this is due to the simple fact that the probability-generating function of $W$ is $g(z)=e^{\nu(z-1)}$, and $\mathbb{E}\left[W^{\underline{m}}\right]=g^{(m)}(1)$. By (15), the $m$-th factorial moment of $C_{n}$ thus becomes
$\mathbb{E}\left[C_{n}^{\underline{m}}\right]=\mathbb{E}\big{[}\mathbb{E}[C_{n}^{\underline{m}}\mid L_{n}]\big{]}=\mathbb{E}[(\mu L_{n})^{m}]=\mu^{m}\mathbb{E}\left[L_{n}^{m}\right],\quad m>0.$ (16)
Together with the identity
[…]
where {${}{m}{i}$ is a Stirling number of the second kind, this gives the moments of $C_{n}$ in terms of the moments of $L_{n}$ as
[…]
For example, the first four moments read
$\begin{split}\mathbb{E}\left[C_{n}\right]&=\mu\mathbb{E}\left[L_{n}\right],\\ \mathbb{E}\left[C_{n}^{2}\right]&=\mu\mathbb{E}\left[L_{n}\right]+\mu^{2}\mathbb{E}\left[L_{n}^{2}\right],\\ \mathbb{E}\left[C_{n}^{3}\right]&=\mu\mathbb{E}\left[L_{n}\right]+3\mu^{2}\mathbb{E}\left[L_{n}^{2}\right]+\mu^{3}\mathbb{E}\left[L_{n}^{3}\right],\quad\text{and}\\ \mathbb{E}\left[C_{n}^{4}\right]&=\mu\mathbb{E}\left[L_{n}\right]+7\mu^{2}\mathbb{E}\left[L_{n}^{2}\right]+6\mu^{3}\mathbb{E}\left[L_{n}^{3}\right]+\mu^{4}\mathbb{E}\left[L_{n}^{4}\right].\end{split}$ (18)
In particular, the variance is
$\begin{split}\mathbb{V}[C_{n}]&=\mathbb{E}\left[C_{n}^{2}\right]-(\mathbb{E}\left[C_{n}\right])^{2}\\ &=\mu\mathbb{E}\left[L_{n}\right]+\mu^{2}(\mathbb{E}\left[L_{n}^{2}\right]-(\mathbb{E}\left[L_{n}\right])^{2})\\ &=\mu\mathbb{E}\left[L_{n}\right]+\mu^{2}\mathbb{V}[L_{n}].\end{split}$ (19)
The latter can be seen as an instance of the standard decomposition of the variance: $\mathbb{V}\left[C_{n}\right]=\mathbb{E}\left[\mathbb{V}\left[C_{n}\mid L_{n}\right]\right]+\mathbb{V}\left[\mathbb{E}\left[C_{n}\mid L_{n}\right]\right]$. So the first term, $\mathbb{E}\left[\mathbb{V}\left[C_{n}\mid L_{n}\right]\right]=\mu\mathbb{E}\left[L_{n}\right]=\mathbb{E}\left[C_{n}\right]$, contains the variability due to the innovation process, whereas the second term, $\mathbb{V}\left[\mathbb{E}\left[C_{n}\mid L_{n}\right]\right]=\mu^{2}\mathbb{V}\left[L_{n}\right]$, comes from the variance of the length of the genealogy.
3.1.2 Path integrals of birth-death processes
Now that we know the moments of $C_{n}$ in terms of the moments of $L_{n}$, we set out to determine the latter. Note that $L_{n}$ in (10) is a special case of the path integral of a function $f:\mathbb{N}_{0}\to\mathbb{R}_{\geq 0}$ over a general continuous-time birth-death process $(Z(t))_{t\geqslant 0}$ on $\mathbb{N}_{0}$ with unique absorbing state 0, namely
$\mathcal{I}_{i}(f):=\int_{0}^{\infty}f(Z(t)\mid Z(0)=i)\mathrm{d}{t},\quad i>0.$ (20)
We assume throughout [^3] Note that, on this finite state space, we need not assume that $f$ is nondecreasing for large arguments, as required for $(Z(t))_{t\geqslant 0}$ on $\mathbb{N}_{\geq 0}$ in Stefanov and Wang (2000). that $f(0)=0$ and $f(i)>0$ for $i\in[N]$; in particular, this implies that the integral has no contribution from beyond the time of absorption in 0. The mathematical features of $\mathcal{I}_{i}(f)$, especially its expectation and its higher-order moments, have been well studied (Puri, 1966, 1968; Jerwood, 1970; McNeil, 1970; Goel and Richter-Dyn, 1974; Norden, 1982; Hernandez-Suarez and Castillo-Chavez, 1999; Stefanov and Wang, 2000; Pollett and Stefanov, 2002; Pollett, 2003; Crawford et al., 2018; Hobolth et al., 2019). Here we restrict $(Z(t))_{t\geqslant 0}$ to $[N]_{0}$. In Appendix B, we outline how to derive the moments of the path integral via first-step analysis; below we summarise the results.
For $j\in[N]_{0}$, let $\lambda_{j}$ and $\mu_{j}$, respectively, be the birth and death rates, that is, the rates for the transitions $j\to j+1$ and $j\to j-1$, in $(Z(t))_{t\geqslant 0}$; we assume that $\lambda_{0}=0$ (since 0 is absorbing), $\lambda_{j}\geq 0$ for $j\in[N-1]$, and $\mu_{j}>0$ for $j\in[N]$, and complement this by the convention $\mu_{0}=\lambda_{N}=0$. The expectation of $\mathcal{I}_{i}(f)$ in (20) is then given by
$\mathbb{E}\left[\mathcal{I}_{i}(f)\right]=\sum_{j=1}^{N}\zeta_{ij}f(j),$ (21)
where, for $i,j\in[N]$,
$\zeta_{ij}\mathrel{\mathop{:}}=\sum_{\ell=1}^{\min\{i,j\}}\eta_{\ell j},$ (22)
and, for $1\leq\ell\leq j\leq N$,
$\eta_{\ell j}\mathrel{\mathop{:}}=\frac{\lambda_{\ell}\cdots\lambda_{j-1}}{\mu_{\ell}\cdots\mu_{j}},$ (23)
where the empty product is 1 (p. 162/163 in Stefanov and Wang (2000); note that, for $\ell\leq j$, $\eta_{\ell j}$ equals their $H_{\ell}(j)/\mu_{\ell}$; see also Hernandez-Suarez and Castillo-Chavez (1999) for the special case of $i=1$, and see eq. (1) of Stefanov and Wang (2000) for the limit $N\to\infty$). Here, $\eta_{\ell j}$ represents the expected total sojourn time in state $j$ measured from the moment where the process reaches state $\ell$ for the first time until the moment where it reaches state $\ell-1$ for the first time; for lack of reference and convenience of the reader, we prove this fact in Appendix A. Likewise, $\zeta_{ij}$ is the expected total sojourn time of $(Z(t))_{t\geqslant 0}$ in state $j$ before the process is absorbed in 0, given it started in $i$ (Stefanov and Wang, 2000, Eq. (4)). Generalising (21), the $m$-th moment of the path integral is given by
$\mathbb{E}\left[\big{(}\mathcal{I}_{i}(f)\big{)}^{m}\right]=m!\sum_{j_{1}=1}^{N}\cdots\sum_{j_{m}=1}^{N}\zeta_{ij_{1}}\zeta_{j_{1}j_{2}}\cdots\zeta_{j_{m-1}j_{m}}f(j_{1})\cdots f(j_{m}),$ (24)
see Appendix B.
3.1.3 Tree length
Let us now apply the results above to our cultural evolution model. With the choice $(Z(t))_{t\geqslant 0}=(Y(\tau))_{\tau\geqslant 0}$, $f=\mathrm{id}$, and $i=n$, we have $L_{n}=\mathcal{I}_{n}(f)$. With the transition rates of $(Y(\tau))_{\tau\geqslant 0}$ in (9), (23) evaluates to
$\eta_{\ell j}=\frac{1}{uj}\left(\frac{s}{Nu}\right)^{j-\ell}\frac{(N-\ell)!}{(N-j)!}=\frac{1}{uj}\left(\frac{s}{Nu}\right)^{j-\ell}(N-\ell)^{\underline{j-\ell}},\quad 1\leq\ell\leq j\leq N,$ (25)
and (22) becomes
$\zeta_{nj}=\sum_{\ell=1}^{\min\{n,j\}}\frac{1}{uj}\left(\frac{s}{Nu}\right)^{j-\ell}(N-\ell)^{\underline{j-\ell}}.$ (26)
In Appendix C, we show that, with the help of (26), (21) and (24) evaluate to
$\mathbb{E}\left[L_{n}\right]=\sum_{m=0}^{N-1}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m}\frac{N^{\underline{m+1}}-(N-n)^{\underline{m+1}}}{m+1}$ (27)
and
$\begin{split}\mathbb{E}\left[L_{n}^{2}\right]&=2\sum_{m_{1}=0}^{N-1}\sum_{m_{2}=0}^{N-1}\left(\frac{1}{u}\right)^{2}\left(\frac{s}{Nu}\right)^{m_{1}+m_{2}}\\ &\qquad\times\left[\frac{\{N^{\underline{m_{1}+1}}-(N-n)^{\underline{m_{1}+1}}\}N^{\underline{m_{2}+1}}}{(m_{1}+1)(m_{2}+1)}-\frac{N^{\underline{m_{1}+m_{2}+2}}-(N-n)^{\underline{m_{1}+m_{2}+2}}}{(m_{1}+m_{2}+2)(m_{2}+1)}\right].\end{split}$ (28)
Inserting these results into eqs.(18) and (19), we obtain the moments of $C_{n}$ in a closed form, see Figure 4 for expectation and variance. This also illustrates the (smooth) transition from the subcritical phase with a small number of traits to the supercritical phase, where the mean number of traits increases steeply with $s$, as previously observed in similar models (Nakamura et al., 2020; Kobayashi et al., 2018; Kobayashi et al., 2021). Comparing with the variance decomposition below (19), the figure also shows that, for $s>1$, the variability due to innovations, which equals $\mathbb{E}\left[C_{N}\right]$, is small relative to $\mathbb{V}\left[C_{N}\right]$, so the major contribution to $\mathbb{V}\left[C_{N}\right]$ is $\mu^{2}\mathbb{V}\left[L_{N}\right]$, which stems from the variance of the tree length. We compare the result to simulations in the next section.
Figure 4: $\mathbb{E}\left[C_{N}\right]$ (solid) and $\mathbb{V}\left[C_{N}\right]$ (broken) as functions of $s$. $N=100,\mu=0.1,u=1$.
3.2 Forward simulations
3.2.1 Procedure
We carried out individual-based simulations of the forward process $(\Phi(t))_{t\in[0,t_{\rm{max}}]}$ described in Section 2.2. Recall that death, learning, and innovation events occur at total rates $Nu$, $Ns$, and $N\mu$, respectively. We work with $u=1$ throughout, so that a time interval of length 1 corresponds to the expected life time of an individual, that is, one generation. The waiting time and kind for each next event is determined by drawing exponential random numbers with parameter $N(u+s+\mu)$ and then letting a death, a learning, and an innovation event happen with probability $Nu/(N(u+s+\mu))$, $Ns/(N(u+s+\mu))$, and $(N\mu/N(u+s+\mu))$, respectively. When a death event occurs, a uniformly-chosen individual loses all traits. When learning occurs, we choose two individuals uniformly and with replacement, namely a teacher (that is, a cultural parent) and a learner, and all traits of the teacher are added to the learner’s repertoire. When an innovation occurs, a uniformly-chosen individual acquires a new trait that has never existed before. Within the limited memory of the computer, this is feasible by "forgetting" the extinct traits and reusing the corresponding storage.
We start with $k_{i}(0)=\varnothing$ for every $i\in[N]$ and compute realisations $c_{n}(t)$ of $C_{n}(t)$ at $t=0,1,...,t_{\rm{max}}$. We take $t_{\rm{max}}$ large enough so that the effect of the initial state can be neglected and denote by $[t_{\rm{max}}]_{0}$ the set $\{0,1,\ldots,t_{\rm{max}}\}$. Note that, by definition (see (5)), $C_{n+1}(t)\geq C_{n}(t)$ for all $t$.
3.2.2 Time series
A cutout of the time series $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ of a subcritical case ($s<u$) and a supercritical case ($s>u$) are shown in Figure 5. More detail is to be seen in Figure 6, which shows the time series $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ for the last 400 generations in both cases. In the subcritical case, both $(c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ fluctuate moderately around some constant value in a more or less symmetric manner. In contrast, we find a characteristic sawtooth behaviour in the supercritical case: $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ tends to increase linearly (at roughly constant slope in every sawtooth) for some period, followed by a rapid collapse to a small value at a random time. We will discuss the sawtooth behaviour later. For now, note that $t_{\rm{max}}$ is large enough for the mass extinction of traits to occur sufficiently often during the simulated period. Despite the dynamics, we are thus in the stationary situation. Let us therefore first consider the moments of (the stationary) $C_{n}$ and then turn to the dynamics.
Figure 5: Time series $\big{(}c_{N}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$). The last $10^{4}$ generations of a long run with $t_{\rm{max}}=10^{5}$ are shown. $N=100,\mu=0.1,u=1.$
Figure 6: Time series $\big{(}c_{1}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ (green) and $\big{(}c_{N}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ (blue) for the last 400 generations in a long run with $t_{\rm{max}}=10^{5}$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$). $N=100,\mu=0.1,u=1$.
3.2.3 Moments of $C_{n}$
As a consistency check, we calculate the empirical time average and the corresponding variance, that is,
$\bar{c}_{n}\mathrel{\mathop{:}}=\frac{1}{t_{\rm{max}}+1}\sum_{i=0}^{t_{\rm{max}}}c_{n}(t)$ (29)
and
$s_{c}^{2}\mathrel{\mathop{:}}=\frac{1}{t_{\rm{max}}}\sum_{i=0}^{t_{\rm{max}}}\big{(}c_{n}(t)-\bar{c}_{n}\big{)}^{2}$ (30)
for various values of $n$ and compare $\bar{c}_{n}$ and $\sqrt{s_{c}^{2}}$ with $\mathbb{E}\left[C_{n}\right]$ and $\sqrt{\mathbb{V}\left[C_{n}\right]}$ according to eqs.(18) and (19). Figure 7 shows that they are in good agreement in both the sub- and supercritical case. The most important difference is that, in the supercritical case, $\mathbb{E}\left[C_{n}\right]$ saturates for relatively small $n$; that is, almost all traits of the population are already contained in a small sample, as one may already suspect from Figure 6, and as will be analysed in what follows. Of course, in addition, $\mathbb{E}\left[C_{n}\right]$ and $\mathbb{V}\left[C_{n}\right]$ are altogether substantially higher in the supercritical case; also compare Figure 4, which shows the dependence on $s$. Note also that, in the supercritical case, $\mathbb{E}\left[C_{n}\right]\ll\mathbb{V}\left[C_{n}\right]$ for all $n$, so the variance of the tree length dominates the variability due to innovations for all $n\in[N]$, as already discussed in the context of Figure 4 for $n=N$.
Figure 7: Analytic and simulation results of mean and standard deviation of $C_{n}$ as a function of $n$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$). Solid and broken lines show $\mathbb{E}\left[C_{n}\right]$ and $\sqrt{\mathbb{V}\left[C_{n}\right]}$ of (18) and (19), respectively. Large and small discs represent $\bar{c}_{n}$ and $\sqrt{s_{c}^{2}}$ of (29) and (30). $N=100,\mu=0.1,u=1,t_{\rm{max}}=10^{5}$.
3.2.4 Marginal and joint distributions of $C_{n}$
Figure 8 shows the histograms of $(\log_{10}c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(\log_{10}c_{N})_{t\in[t_{\rm{max}}]_{0}}$ in the sub- and supercritical cases. In contrast to the subcritical case, the histogram of $(\log_{10}c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ in the supercritical case is bimodal. It peaks sharply at 0, where most of the mass is located, decays quickly from there, and has a second peak around $c_{1}=10^{2.3}$. So, an individual is ignorant most of the time, but occasionally becomes very knowledgeable. The histogram of $(\log_{10}c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ has a single peak near the second peak of the histogram of $(\log_{10}c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$. What this means becomes clear in the heat maps (Figure 9), which show how the $c_{n}$ covary. In the subcritical case, the correlation between $c_{1}$ and $c_{N}$ is altogether weak; $c_{1}$ is bimodal in the rare case that $c_{N}$ is large. A similar tendency exists between $c_{2}$ and $c_{N}$. In contrast, the correlation is very strong in the supercritical case. When $c_{N}$ is not too small, the distribution of $c_{1}$ is bimodal, meaning that there are two extreme types of individuals: those who know nearly nothing and those who have almost all traits currently present in the population, with very few intermediate cases. An analogous observation applies to $c_{2}$.
Figure 8: Histograms of $\big{(}\log_{10}c_{1}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ (green) and $\big{(}\log_{10}c_{N}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ (violet) obtained by simulation runs of the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$); the overlap area of the two histograms appears blue-green. $c_{1}=0$ was observed 76376 times (upper panel) and 64394 times (lower panel), which are not shown. $N=100,\mu=0.1,u=1,t_{\rm{max}}=10^{5}$.
Figure 9: Heat maps for $(c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ and for $(c_{2}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$) in a long run with $t_{\rm{max}}=10^{5}$. The angle bisector represents $c_{1}(t)=c_{N}(t)$ for all $t$, that is, individual 1 has all traits present in the population at time $t$. $N=100,\mu=0.1,u=1$.
3.3 Genealogies and backward simulations
3.3.1 Distributions of tree length, $L_{n}$
To understand the behaviour observed in the forward simulations, we now recall that $C_{n}\sim\text{Poi}(\mu L_{n})$ together with the observation that, in the supercritical case, the variance of $C_{n}$ mainly stems from the variance of $L_{n}$. We therefore expect that the underlying genealogies play the decisive role and so turn to them now. Figure 10 shows the histograms of $\log_{10}\ell_{1}$ and $\log_{10}\ell_{N}$ (where $\ell_{n}$ is the realisation of the (stationary) $L_{n}$) obtained via (10) by simulating the untyped ALG $(\Lambda(\tau))_{\tau\geq 0}$ according to (6b) until extinction, a large number of times. (Since we need the full untyped ALG in what follows, we refrained from simulating the more efficient line-counting process $(Y_{n}(\tau))_{\tau\geq 0}$ instead.) Note that the left peaks of the histograms of $\log_{10}\ell_{1}$ translate into the left peaks of the histograms of $(\log_{10}c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ in that short tree lengths mean high chances for no or only a small number of innovation events, resulting in (almost) ignorant individuals; and vice versa for the other peaks.
Figure 10: Histogram of $\log_{10}\ell_{1}$ (green) and $\log_{10}\ell_{N}$ (blue), obtained from $10^{4}$ simulation runs of the untyped ALG; the overlap area of the two histograms appears blue-green. In each run, the ancestors were traced until they disappeared. Top: $s=0.9$. Bottom: $s=1.3$. $N=100,u=1$.
3.3.2 Metastability and coalescence in the ALG
Still, we are at the descriptive level. For a true understanding of the behaviour, we must consider the dynamics of the ALG. The behaviour of the line-counting process $(Y_{n}(\tau))_{\tau\geq 0}$ is very well studied, since, as mentioned above, it is, at the same time, the stochastic logistic process of the SIS model in epidemiology. Recall that $(Y_{n}(\tau))_{\tau\geq 0}$ dies out with probability 1; but it is decisive what happens before extinction. In the subcritical case, $(Y_{n}(\tau))_{\tau\geq 0}$ dies out quickly almost surely. In contrast, in the supercritical regime, it is well known (Andersson and Djehiche, 1998; Foxall, 2021) to have a metastable state around $N\bar{\xi}$, where $\bar{\xi}$ is the stable equilibrium of the differential equation (14). Note that $\lambda_{\lfloor N\bar{\xi}\rfloor}\approx\mu_{\lfloor N\bar{\xi}\rfloor}$; also recall from Section 2.6 that the qualitatively different behaviour in the sub- and supercritical regimes reflects the transcritical bifurcation of the equilibria of the differential equation.
From now on, we mainly focus on the supercritical case in order to understand the sawtooth behaviour for $s>u$. For $(Y_{1}(t))_{\tau\geq 0}$ in the supercritical case, there is the following dichotomy: with probability $u/s$, $(Y_{1}(\tau))_{\tau\geq 0}$ goes extinct quickly; otherwise, it grows to reach the metastable state (say $\lfloor N\bar{\xi}\rfloor$ for definiteness) in a short time. $(Y_{N}(\tau))_{\tau\geq 0}$ always moves to $\lfloor N\bar{\xi}\rfloor$ quickly. By a crude approximation, the expected first-passage times from 1 to $\lfloor N\bar{\xi}\rfloor$ (conditional on non-extinction before reaching $\lfloor N\bar{\xi}\rfloor$), and from $N$ to $\lfloor N\bar{\xi}\rfloor$, are both bounded by $N^{2}/u$ (Andersson and Djehiche, 1998, Proof of Lemma 3) [^4] The factor of $1/u$ comes from the fact that Andersson and Djehiche (1998) work with $u=1$.. (According to Andersson and Djehiche (1998), it can be shown with the help of refined arguments that the bound is actually $\mathcal{O}(\log(N))$.)
Whenever the process has reached the metastable state, it fluctuates around it for a long time with fluctuations of order $\sqrt{N}$, before finally going extinct in a rare and rather sudden event. See the comprehensive work of Foxall (2021) for the details. The expectation of the extinction time $T$ when starting from the metastable state, or, more generally, from any initial value $\mathcal{O}(N)$, in a population of size $N$ is given by
$\mathbb{E}\left[T\right]=\frac{s}{(s-u)^{2}}\sqrt{\frac{2\pi}{N}}e^{N\left\{\log(s/u)+(u/s)-1\right\}}(1+{\scriptstyle\mathcal{O}}(1)),$ (31)
and the distribution of $T/\mathbb{E}\left[T\right]$ converges to an exponential distribution with parameter 1, see Andersson and Djehiche (1998); Doering et al. (2005); Nåsell (2011, eq. (12.2)), and Foxall (2021). Since $\log(x)+1/x\geq 1$ for $x\geq 1$ with equality if and only if $x=1$ (note that $\log(1)+(1/1)-1=0$ and $(\,\mathrm{d}/\,\mathrm{d}x)(\log(x)+(1/x)-1)=(1/x)-(1/x^{2})>0$ for $x>1$), (31) means that the time the process spends in the metastable regime increases exponentially with $N$. Our previous attributes such as ‘quickly’ and ‘in a short time’ are actually meant relative to this exponential time scale.
But the line-counting process alone does not suffice to understand what is really going on; rather, one has to consider the dynamics of the ALG, more precisely, the processes $(\Lambda_{\alpha}(\tau))_{\tau\geq 0}$, $\alpha\in[N]$, and $(\Lambda_{[N]}(\tau))_{\tau\geq 0}$, simulated according to (6b). In line with the dichotomy described above, we observe in Figure 11 that the ancestors of a single individual either die out quickly or grow to a metastable number and survive for a long time; likewise, the number of ancestors of the entire population reaches the metastable regime quickly. But we also see that, soon after both $(\Lambda_{\alpha}(\tau))_{\tau\geq 0}$ and $(\Lambda_{[N]}(\tau))_{\tau\geq 0}$ have reached a metastable size, they become identical. This is simply due to the fact that $\Lambda_{\alpha}(\tau)\subseteq\Lambda_{[N]}(\tau)$ for all $\tau\geq 0$, and the sizes of the two sets perform fluctuations around $N\bar{\xi}$ of order $\mathcal{O}(1/\sqrt{N})$, in the course of which they finally meet and hence become identical. From this point onwards, we have $\Lambda_{\alpha}(\tau)=\Lambda_{[N]}(\tau)$ due to (7).
The coalescence events [^5] Note that these coalescence events are different from those in the ASG, where coalescence means that two lines merge into one. responsible for this happen whenever, for some $\alpha$, a line in $\Lambda_{\alpha}(\tau)$ learns from a line in $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$: then, a parent from $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$ is added to $\Lambda_{\alpha}(\tau)$ and hence removed from $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$. A transition in the reverse direction (that is, an addition of elements to $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$) is not possible: when a line in $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$ learns from a line in $\Lambda_{\alpha}(\tau)$, neither $\Lambda_{\alpha}(\tau)$ nor $\Lambda_{[N]}(\tau)$ will change.
Figure 11: Simulation of the untyped ALG. $\lvert(\Lambda_{[N]}(\tau))_{\tau\geq 0}\rvert$, the number of ancestors of the total population over time (thick grey line), and the $\lvert(\Lambda_{\alpha}(\tau))_{\tau\geq 0}\rvert$, the sizes of the ancestries of each individual $\alpha$ (100 thin black lines for $\alpha\in[N]$) are shown from backward time $\tau$=0 to 40, all based on the same realisation of $(\Lambda(\tau))_{\tau\geq 0}$. By definition, the thick grey line starts at $N$, and each black line starts at 1. When a black line merges into the grey line (26 occurrences among 100, which is close to the probability of $0.231-u/s=$ 0.23), their ancestor sets become identical and behave identically from then on. $N=100,s=1.3,u=1$.
Nevertheless, this process is not one-dimensional: there are also death events in $\Lambda_{\alpha}(\tau)$ and $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$, as well as learning events from parents outside $\Lambda_{N}(\tau)$, by which either of the two sets increases or decreases individually (these types of events are responsible for the fluctuations around the metastable state).
All this carries over to the traits and explains the stationary behaviour observed in Section 3.2.4. In the subcritical case, the ancestry of all individuals dies out quickly without much of a chance to coalesce, so they only acquire few traits in a more or less independent way; as a consequence, $C_{1}$ and $C_{N}$ both remain small with little dependence between them. This changes in the supercritical case. If the ancestry of a given individual dies out quickly, it will again acquire no or only a few traits; but if it has old ancestors, it will accumulate numerous traits. Moreover, due to (relatively fast) coalescence, the set of old ancestors, and thus the set of traits, is largely shared with the entire population. This explains the bimodal behaviour, that is, $L_{1}$ and $C_{1}$ are either very small or close to $L_{N}$ and $C_{N}$, respectively. Put differently, cultural traits that are old but not too old to have gone extinct are carried by all knowledgeable individuals. So, for old traits that still exist in a population, there are two main types of individuals: those who know none of them and those who know all of them. Altogether, the trait diversity between individuals is low.
3.3.3 Evolving genealogies
So far, we have understood the stationary behaviour on the basis of $C_{n}\sim\text{Poi}(\mu L_{n})$. It remains to understand the dynamics, that is, the sawtooth behaviour in the supercritical case. To this end, and once more on the basis of the Poissonian relationship between the number of traits and the tree length, we now consider the tree length as a function of forward time $t$ in a stationary version of the forward process, coupled across all times, and define it as
$L_{n}^{t}\mathrel{\mathop{:}}=\int_{0}^{t}\lvert\Lambda_{[n]}^{t}(\tau)\rvert\mathrm{d}\tau=\int_{0}^{T^{t}}\lvert\Lambda_{[n]}^{t}(\tau)\rvert\mathrm{d}\tau,\quad n\in[N],$ (32)
where $\Lambda_{[n]}^{t}(\tau)$ is the set of ancestors at backward time $\tau$ of individuals $1,2,\ldots,n$ at forward time $t$, see Figure 12 for an illustration; and $T^{t}$ is the extinction time of $(\Lambda^{t}_{[N]}(\tau))_{\tau\geq 0}$, which we take to be equal to $t$ if the process has not died out until backward time $t$. We clearly have
$\Lambda^{t}_{[n]}(0)=[n].$ (33)
As before, the set $[n]$ is representative of any sample of size $n$ due to exchangeability. Let us also note that the $T^{t}$ are all identically distributed in the same way as $T$ of (31), but they are, in general, not independent; and likewise for $L_{n}^{t}$ and $\Lambda_{[n]}^{t}$. All this leads us to the concept of evolving genealogies, as previously studied in the context of the neutral coalescent process (that is, without selection) by Pfaffelhuber et al. (2011). The time course of $(C_{n}(t))_{t\geq 0}$ will follow from there via $C_{n}(t)\sim\textrm{Poi}(\mu L_{n}^{t})$.
Figure 12: The total branch length $L_{N}^{t}$ of the entire population as a function of forward time $t$. The realisation of the forward model, as given by the positions and types of all events, is identical in both panels. The left (right) panel shows a case with $0<t<t_{\rm{max}}$ ($t=t_{\rm{max}}$). The sum of the lengths of the red line segments is $L_{N}^{t}$.
We have extracted the evolving genealogies from a single long simulation run of the forward model in the supercritical case, where we have stored all events (including the information about who died and who learned from whom) and the times at which they happen. This way, we obtained the realisations of $(L^{t}_{N})_{t\geq 0}$ and $(L^{t}_{1})_{t\geq 0}$ coupled across all times, as shown in Figure 13.
Figure 13: The tree length (red) and the number of traits (blue) as functions of forward time $t$ in the same realisation of a forward simulation. Top: $\mu(\ell_{N}^{t})_{t\geq 0}$ and $(c_{N}(t))_{t\geq 0}$. Bottom: $\mu(\ell_{1}^{t})_{t\geq 0}$ and $(c_{1}(t))_{t\geq 0}$. In the bottom panel, they both oscillate quickly between near-0 and close to $\mu\ell_{N}^{t}$, so only points (instead of lines) are shown. $N=100,s=1.3,\mu=0.1,u=1$.
Figure 14: Relationship between (from top to bottom) the total tree length, the set of ancestors, the time to the oldest ancestor, and the descendant process, all extracted from the same forward simulation, and all as a function of forward time $t$. In detail: the top panel shows a realisation $(\ell_{N}^{t})_{t\geq 0}$ of $(L_{N}^{t})_{t\geq 0}$ with a mass extinction at $t\approx 9688$. The second panel displays the number of ancestors at $t\leq t_{i}$ ($i=1,2,3,4)$ of samples taken at forward times $t_{1}=9681,t_{2}=9706,t_{3}=9748,t_{4}=9789$. More precisely, the thick coloured lines are realisations of the $\lvert\Lambda^{t_{i}}_{[N]}(t_{i}-t)\rvert_{t\leq t_{i}}$, whereas the thin grey lines represent realisations of $\lvert\Lambda^{t_{i}}_{\alpha}(t_{i}-t)\rvert_{t\leq t_{i}}$ for all $\alpha\in[N]$; these lines get darker when more lines overlap. The third panel shows the realisation $(s_{N}(t))_{t\geq 0}$ of $(S_{N}^{t})_{t\geq 0}$, the time to the oldest ancestor. The bottom panel displays the number of descendants at time $t\geq t_{i}$ of samples taken at the forward times $t_{i}$. In analogy with the second panel, the thick coloured lines and the thin grey lines are realisations of the $\lvert\Gamma^{t_{i}}_{[N]}(t-t_{i})\rvert_{t\geq t_{i}}$ and the $\lvert\Gamma^{t_{i}}_{\alpha}(t-t_{i})\rvert_{t\geq t_{i}}$, respectively. Looking backward (ancestral process) or forward (descendant process) in time, the thick coloured lines (along with the thin black ones that have coalesced into them) merge and fluctuate around their metastable values for a long time before going extinct. $N=100,u=1,s=1.3.$
Indeed, the realisation shows the sawtooth picture with linear increase of $(L_{N}^{t})_{t\geq 0}$ interrupted by sudden near-extinctions at random times. But to understand it, we need the finer picture of $\lvert\Lambda^{t}_{[n]}(\tau)\rvert_{\tau\geq 0}$ in Figure 14. The figure shows realisations of $\lvert\Lambda^{t_{i}}_{[n]}(t_{i}-t)\rvert_{t\leq t_{i}}$ for $i\in\{1,2,3,4\}$, $n\in\{1,N\}$, and starting times $t_{i}$ chosen so that
$t_{1}\lesssim t_{2}^{\rm e}=t_{3}^{\rm e}=t_{4}^{\rm e}\ll t_{2}\ll t_{3}\ll t_{4},$ (34)
where $t_{i}^{\rm e}$ is the time point where $(\Lambda^{t_{i}}_{[N]}(t_{i}-t)_{t\leq t_{i}})$ goes extinct; so $t_{i}-t_{i}^{\rm e}$ is the corresponding realisation of $T^{t_{i}}$. Note that, due to (34), the $t_{i}-t_{i}^{\rm e}$ are coupled for $i\in\{2,3,4\}$. With $\lesssim$ and $\ll$, we indicate that the quantities are close and not close to each other, respectively. We now see that the realisations of $(\Lambda^{t_{3}}_{[N]}(t_{3}-t))_{t\leq t_{3}}$ and $(\Lambda^{t_{2}}_{[N]}(t_{2}-t))_{t\leq t_{2}}$ quickly get close to, and actually become identical with, the realisation of $(\Lambda^{t_{4}}_{[N]}(t_{4}-t))_{t\leq t_{4}}$; in particular, they join into the same metastable set and are extinguished with it. Likewise, the realisations of $(\Lambda^{t_{i}}_{\alpha}(t_{i}-t))_{t\leq t_{i}}$, $i\in\{2,3,4\}$ and $\alpha\in[N]$, either die out quickly or coalesce with the metastable set. This coalescence is clear by arguments similar to those in Section 3.3.2: since $\Lambda^{t_{4}}_{[N]}(t_{4}-t_{3})\subseteq\Lambda^{t_{3}}_{[N]}(0)=[N]$, we have $\Lambda^{t_{4}}_{[N]}(t_{4}-t_{3}+\tau)\subseteq\Lambda^{t_{3}}_{[N]}(\tau)$ for all $\tau>0$; so, by moving towards the metastable size, the latter process joins into the former from above. Likewise, $\Lambda^{t_{3}}_{\alpha}(0)\subseteq\Lambda^{t_{3}}_{[N]}(0)$, so $\Lambda^{t_{3}}_{\alpha}(\tau)\subseteq\Lambda^{t_{3}}_{[N]}(\tau)$ for $\tau>0$; hence the former process, if it does not die out quickly, joins into the latter from below. Analogous arguments hold for the other time points.
Now, since $\lvert\Lambda^{t}_{[N]}(\tau)\rvert_{\tau\geq 0}$ spends most of its time alive near $N\bar{\xi}$, (32) tells us that $L_{N}^{t}\approx N\bar{\xi}T^{t}$. In our specific realisation, we have $\ell_{N}^{t}\approx N\bar{\xi}(t-t_{4}^{\rm e})$ for any time $t_{4}^{\rm e}\lesssim t\lesssim t_{4}$ due to the identity of the $t_{i}^{\rm e}$, so we see a linear decrease (at rate $\approx N\bar{\xi}=N(1-u/s)$) with decreasing $t$. Restarting at $t_{1}$, in contrast, leads to a new metastable state, which awaits its extinction at $t_{1}^{\rm e}$, independently of $t_{2}^{\rm e}$. Moreover, since $t_{1}$ is close to $t_{2}^{\rm e}$, the value of $\ell_{N}^{t_{1}}$ is close to the maximum of the previous sawtooth. More generally, we are led to conjecture that $L_{N}^{t}$ decreases approximately linearly at rate $N\bar{\xi}$ for decreasing $t$ as long as $T^{t}\gg 0$ and then, after surpassing some small value, moves quickly to a new peak.
So far, we have followed the ancestral process backward in time. But the figure can also be read forward. Let $\Gamma_{[n]}^{t}(\sigma)$ be the set of descendants at time $t+\sigma$, $\sigma\geq 0$, of the set $[n]$ of individuals at forward time $t$. The process $(\Gamma_{[n]}^{t}(\sigma))_{\sigma\geq 0}$ has the same law as $(\Lambda_{[n]}^{t}(\tau))_{\tau\geq 0}$ with $\tau$ replaced by $\sigma$; this follows immediately from the fact that the set of descendants (ancestors) is obtained by following all learning arrows in the forward (backward) direction (and pruning lines that meet a cross), and a reversal of all arrows does not change the law of the process. As a side remark, let us note that $\lvert\Gamma_{[n]}^{0}(t)\rvert_{t\geq 0}$ has the same law as $(X(t))_{t\geq 0}$ of our single-trait model with $X(0)=n$; this reflects the self duality of the SIS model. In any case, like $\lvert\Lambda_{[N]}^{t}(\tau)\rvert_{\tau\geq 0}$, also $\lvert\Gamma_{[N]}^{t}(\sigma)\rvert_{\sigma\geq 0}$ moves to the metastable state around $N\bar{\xi}$ quickly, and like $\lvert\Lambda_{\alpha}^{t}(\tau)\rvert_{\tau\geq 0}$, also $\lvert\Gamma_{\alpha}^{t}(\sigma)\rvert_{\sigma\geq 0}$ dies out quickly with probability $u/s$ and otherwise moves to the metastable state.
The important point now is that the two processes are connected via
[…]
because if an individual $\alpha$ is ancestral to $\beta$, then $\beta$ is a descendant of $\alpha$. In particular, we have
$\Lambda_{[N]}^{t}(t-s)=\{\alpha:\Gamma_{\alpha}^{s}(t-s)\neq\varnothing\}.$ (35)
In general, $\Lambda_{[N]}^{t}(t-s)\neq\Gamma_{[N]}^{s}(t-s)$. However, it is true that
$\Lambda_{[N]}^{t}(t-s)=\varnothing\Longleftrightarrow\Gamma_{[N]}^{s}(t-s)=\varnothing$ (36)
for any $0<s<t$, because if the population at time $t$ does not have ancestors at time $s$, then the descendants of the population at time $s$ do not survive until $t$. Moreover, if $t\gg s$ and the sets in (36) are not empty, we have
$\lvert\Lambda_{[N]}^{t}(t-s)\rvert\approx N\bar{\xi}\approx\lvert\Gamma_{[N]}^{s}(t-s)\rvert$ (37)
due to metastability. In the realisation of the figure, the equalities in (36) are true for $s=t_{1}$, $t=t_{2}$; that is, both processes go extinct in $[t_{1},t_{2}]$. In contrast, for $s=t_{2}$ and $t=t_{3}$ or $t=t_{4}$, (37) applies.
Now, (36) and (37) together allow us to rewrite $L_{N}^{t}$ of (32) as
$L_{N}^{t}=\int_{0}^{t}\lvert\Gamma_{[N]}^{s}(t-s)\rvert\,\mathrm{d}s=\int_{0}^{S^{t}}\lvert\Gamma_{[N]}^{t-\sigma}(\sigma)\rvert\,\mathrm{d}\sigma\approx S^{t}N\bar{\xi},$ (38)
where $S^{t}$ denotes the largest $\sigma\leq t$ such that $\Gamma_{[N]}^{t-\sigma}(\sigma)\neq\varnothing$, or, equivalently, the smallest $\sigma\leq t$ such that $\Gamma_{[N]}^{(t-\sigma)-}(\sigma)=\varnothing$. This is the time [^6] Note that $(\Gamma_{[n]}^{t}(\sigma))_{\sigma\geq 0}$ is based on the forward process, so it is càdlàg; that is, a jump at time $t$ means that the ’old’ state applies until time $t-$ (the moment ‘just before’ $t$), and at time $t$, the process is already in the ‘new’ state. This implies that the ancestors are extinct at $(t-S^{t})-$, but alive at $t-S^{t}$. since the last extinction event before $t$ (assumed to be $t$ if no extinction has taken place before $t$). In other words, the population at forward time $t$ has their oldest ancestor(s) at forward time $t-S^{t}$. Since the extinction times are distinct random points on the time axis, $(S^{t})_{t\geq 0}$ is a sawtooth function that increases linearly with slope 1 and is reset to some small value at random times. So (38) explains the linear increase in forward time, interrupted by steep descents at the jumps of $(S_{t})_{t\geq 0}$.
We can now also understand the dynamics of $(L_{1}^{t})_{t\geq 0}$ as shown in the lower panel of Figure 13. Since, for any given $t$, $(\Lambda^{t}_{\alpha}(\tau))_{\tau\geq 0}$ quickly either dies out or gets close to $(\Lambda^{t}_{N}(\tau))_{\tau\geq 0}$, it is clear that, most of the time, $(L_{1}^{t})_{t\geq 0}$ is either close to 0 or close to $(L_{N}^{t})_{t\geq 0}$. The frequent, random transitions between the two possibilities come from the fact that, at the metastable state over time, the actual set of lines that has old ancestors, while having approximately constant size, moves around in the population via the events of teaching individuals outside the current set, and by death events, thus rapidly including or excluding individuals. That is, by learning from someone with old ancestors, an individual acquires the entire ancestry of the parent; and when the individual dies, it loses all its ancestry.
Let us finally get back to the traits via the relationship $C_{n}(t)\sim\text{Poi}(\mu L_{n}(t))$ and the observation that the variability in $C_{n}$ is largely governed by the variability of $L_{n}$ rather than that of the innovation process. It is then no surprise that the time series $(\mu\ell_{N}^{t})_{t\geq 0}$ and $(c_{N}(t))_{t\geq 0}$ agree almost perfectly in our simulations, and likewise for $(\mu\ell_{1}^{t})_{t\geq 0}$ and $(c_{1}(t))_{t\geq 0}$, as also illustrated in Figure 13. This confirms that the dynamics of the traits results from the genealogical features of the model. In particular, traits are collected at a constant rate while the genealogy is in a metastable state, and most traits are lost simultaneously when the genealogy collapses.
Let us note in passing that, for the parameter values in Figure 13, (31) gives $\mathbb{E}\left[T\right]\approx 85.3$. Manual counting in Figure 13 yields $\approx 83$ sawteeth, which amounts to a mean tooth length (or extinction time) of 10000/83 $\approx$ 120. The overestimation of the length is presumably due to the underestimation of the number of teeth because small teeth are not resolved.
3.4 Popularity spectrum
Aoki (2018) (see also Strimling et al. (2009); Fogarty et al. (2015, 2017)) studied the popularity spectrum $({P_{l}})_{l\in[N]}$, where $P_{l}\mathrel{\mathop{:}}=\mathbb{E}\left[C_{N,l}\right]$ and $C_{n,l}$ is the number of traits carried by exactly $l$ individuals in a sample of size $n$ at stationarity. The aforementioned papers obtained the popularity spectrum in a discrete-time model via the equilibrium condition in the forward model. In Appendix D, we use the corresponding continuous-time version to derive
${P_{l}}=\left(\frac{s}{Nu}\right)^{l-1}\frac{\mu N^{\underline{l}}}{ul},\quad\ell\in[N],$ (39)
in agreement with the simulations in Figure 15. Starting from the $P_{l}$, we also extend the trait frequency spectrum to a sample of size $n$ and obtain in Appendix D that
$\mathbb{E}\left[C_{n,i}\right]=\sum_{l=1}^{N}P_{l}\frac{\binom{l}{i}\binom{N-l}{n-i}}{\binom{N}{n}},\quad i\in[n].$ (40)
As a consistency check, or as an alternative derivation of $\mathbb{E}\left[C_{n}\right]$ via a forward approach, we also get $\mathbb{E}\left[C_{n}\right]=\sum_{i=1}^{n}\mathbb{E}\left[C_{n,i}\right]=\mu\mathbb{E}\left[L_{n}\right]$ of (18) and (27) (note that $\mathbb{E}\left[C_{N}\right]$ is called $C_{\rm{pop}}$ in some of the aforementioned studies).
The bimodal distribution of the popularity spectrum in the supercritical case observed in Figure 15 reflects the by-now-familiar fact that some traits die out quickly while the others survive for a long time, with very few intermediate traits, recall Figure 9.
Figure 15: Popularity spectrum obatained from a long forward simulation run (bullets) compared with the analytic formula (39) (solid line). $s=1.3,N=100,\mu=0.1,u=1,t_{\rm{max}}=10^{5}$.
4 Discussion
Mathematical models for the accumulation of cultural traits in a population over time are crucial for understanding the cultural variation in human populations. Here we have investigated cultural dynamics in which selectively neutral, discrete cultural traits are invented by individuals independently of one another at a constant rate and transmitted between individuals through random social learning. Related previous analyses (Strimling et al., 2009; Lehmann et al., 2011; Fogarty et al., 2015, 2017; Aoki, 2018) were mostly static and centered around the traits, focussing on the expectation of quantities like the number of distinct cultural traits maintained in a population at stationarity. Our approach is different in three ways: first, we concentrate on the underlying genealogies (as first described by Aguilar and Ghirlanda (2015)) and thus obtain results that remain true independently of the traits; and second, we investigate the dynamical aspects, which seem to be unexplored so far. At the heart of this are the time evolution of the set of ancestors of a single individual or the entire population; the coalescent process between the various ancestral sets and the metastability of the corresponding counting processes, which agree with the well-studied stochastic logistic model (Nåsell, 2011; Foxall, 2021); and the concept of evolving genealogies. Third, we also obtain moments of $L_{n}$ and $C_{n}$.
The most conspicuous feature of our model is the sawtooth-like behaviour of $(L_{N}^{t})_{t\geq 0}$ in the supercritical case, which is reflected in $(C_{N}(t))_{t\geq 0}$. The former is an inherent (and, to the best of our knowledge, novel) property of the evolving genealogy and deserves further mathematical investigation. In contrast, the latter is a consequence of the transmission model assumed here for simplicity: in every learning event, all traits carried by the role model are transmitted to the learner, and the traits, once learnt, are never forgotten until death strikes the carrier. Thus, traits are never detached from one another once they come together in an individual. This “sticky” nature of traits implies that (nearly) ignorant individuals immediately become knowledgeable once they learn from a knowledgeable parent. Despite the somewhat artificial assumption, the model may capture certain aspects of the accumulation of technologies or knowledge in human populations. For example, techniques for traditional craftwork are usually products of the accumulation of improvements over many generations and are transmitted as clusters through close apprenticeship; but they can be irreversibly lost if all the carriers happen to die or no successors are found.
Indeed, it is increasingly recognised based on evidence as well as theory that cumulative cultural evolution of human technologies does not occur in a monotonic manner, but via phases of gradual accumulation of innovations punctuated by sudden changes like rapid cascades of innovations (“leaps”) or drastic loss (Kolodony et al., 2015; Vidiella et al., 2022). In particular, the sudden loss of sophisticated or complex technologies and subsequent replacement by degraded ones in ethnographic or archaeological records is often attributed to demographic factors such as population bottlenecks or the fragmentation of social networks, see, for example, Henrich (2004) or Jacobs and Roberts (2009). Our results demonstrate that such behaviour is, in principle, also possible in the absence of demographic changes, via the inherent stochastic properties of the evolving genealogies.
As an outlook, let us nevertheless extend the model to allow for independent transmission of traits, similar to the discrete-time model of Kobayashi et al. (2018), by assuming that each trait of the role model is independently transmitted to the learner with probability $b$ ($b=1$ reproduces our original model).
To keep the mean number of traits transmitted via learning events constant, we replace $s$ by $s/b$. For $b=0.99$ and, hence, our usual choice $s=1.3$ replaced by $s=130/99$, $(c_{N}(t))_{t\geq 0}$ still follows $\mu(\ell_{N}^{t})_{t\geq 0}$ closely, see Figure 16. For $b=13/14$ and, hence, $s=1.3$ replaced by $s=1.4$, the average size of the sawteeth in $(\ell_{N}^{t})_{t\geq 0}$ is significantly higher, in line with the steep increase of $\mathbb{E}\left[L_{N}\right]$ with $s$, see Figure 4. More importantly, the dynamics of $(c_{N}(t))_{t\geq 0}$ does not mirror that of $(\ell_{N}^{t})_{t\geq 0}$ any more. More precisely, a collapse of $(\ell_{N}^{t})_{t\geq 0}$ still implies a collapse of $(c_{N}(t))_{t\geq 0}$, simply because a tiny ancestral tree provides no opportunity to accumulate innovations, no matter how they are transmitted. However, $(c_{N}(t))_{t\geq 0}$ remains below $(\mu\ell_{N}^{t})_{t\geq 0}$, because inheritance of traits is increasingly diluted with their age.
Figure 16: Simulated time series $(\mu\ell_{N}^{t})_{t\in[t_{\rm{max}}]_{0}}$ (red) and $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ (blue) with $b<1$. Upper panel: $s=130/99,b=99/100$. Lower panel: $s=14/10,b=13/14$. Common parameters: $N=100,\mu=0.1,u=1,t_{\rm{max}}=10000$. Where no red line is drawn, it exceeds 700.
Acknowledgment
J.Y.W. received funding from MEXT/JSPS KAKENHI Grant Numbers JP21K03357 and JP24H00001. H.O. acknowledges the support from the "Evolutionary Studies of Complex Adaptive Systems" Research Grant from the Research Center for Integrative Evolutionary Science, SOKENDAI. Y. K. received funding from MEXT/JSPS KAKENHI Grant Numbers 25K06712 and 25K00199. E.B. received funding from the Deutsche Forschungsgemeinschaft, Germany (DFG, German Research Foundation) — Project-ID 317210226 — SFB 1283 and thanks Tom Britton (Stockholm) for an enlightening discussion about the connection to epidemiology.
Appendix
A Sojourn times of the birth-death process
Consider our general continuous-time birth-death process $(Z(t))_{t\geqslant 0}$ on $[N]_{0}$ with unique absorbing state 0 and birth and death rates $\lambda_{j}$ and $\mu_{j}$ with $\lambda_{0}=0$, $\lambda_{j}\geq 0$ for $j\in[N-1]$, and $\mu_{j}>0$ for $j\in[N]$, complemented by $\mu_{0}=\lambda_{N}=0$.
Proposition A.1.
Let $Z(0)=i\in[N]$, let $T_{\ell}$ be the time where the process hits state $\ell\;(1\leq\ell\leq i)$ for the first time, and let $S_{j}^{(\ell)}$ be the total sojourn time in state $j\;(\ell\leq j\leq N)$ in the interval $[T_{\ell},T_{\ell-1}]$. We then have
$\mathbb{E}\big{[}S_{j}^{(\ell)}\mid Z(0)=i\big{]}=\frac{\lambda_{\ell}\cdots\lambda_{j-1}}{\mu_{\ell}\cdots\mu_{j}}=\mathrel{\mathop{:}}\eta_{\ell j},$
where the empty product is 1.
Proof.
The hitting times are finite almost surely due to the almost sure absorption in 0. Fix $j\in[N]$. We first consider $\ell=j$ and let $M_{j}$ be the number of times the chain visits $j$ before moving to $j-1$ (where the initial visit is also counted). When in $j$, the chain moves up and down with probability $\lambda_{j}/(\lambda_{j}+\mu_{j})$ and $\mu_{j}/(\lambda_{j}+\mu_{j})$, respectively, so $M_{j}$ equals $1$ plus a geometric random variable with parameter $\mu_{j}/(\lambda_{j}+\mu_{j})$, where a geometric random variable with success parameter $p$ refers to the number of trials before the first success; note that, for $j=N$, we have the limiting case of $p=1$ and hence $M_{N}=1$. So $\mathbb{E}\left[M_{j}\right]=1+\lambda_{j}/\mu_{j}=(\lambda_{j}+\mu_{j})/\mu_{j}$. Each such visit has a mean duration of $1/(\lambda_{j}+\mu_{j})$, so, by Wald’s identity,
$\mathbb{E}\left[(S_{j}^{(j)}\mid Z(0)=i\right]=\mathbb{E}\left[M_{j}\right]\frac{1}{\lambda_{j}+\mu_{j}}=\frac{1}{\mu_{j}}=\eta_{jj}.$ (A.1)
This proves the claim for $\ell=j$ and serves as the start for induction in $\ell$. For the step from $\ell$ to $\ell-1$, let $N_{\ell-1\to\ell}$ be the number of transitions from $\ell-1$ to $\ell$ in $[T_{\ell-1},T_{\ell-2}]$; then $N_{\ell-1\to\ell}$ follows the geometric distribution with parameter $\mu_{\ell-1}/(\mu_{\ell-1}+\lambda_{\ell-1})$, so $\mathbb{E}\left[N_{\ell-1\to\ell}\right]=\lambda_{\ell-1}/\mu_{\ell-1}$. Let further $S^{(\ell)}_{jn}$ be the sojourn time in $j$ in the interval
[time of the $[\text{time of the }n\text{th arrival in }\ell\text{ after }T_{\ell-1},\text{time of the following arrival in }\ell-1]$
for $n\in[N_{\ell-1\to\ell}]$; clearly, $S^{(\ell)}_{j1}$ equals $S^{(\ell)}_{j}$ in distribution, and, by the Markov property, the $S^{(\ell)}_{jn}$ are iid. Since $S^{(\ell-1)}_{j}=\sum^{N_{\ell-1\to\ell}}_{n=1}S^{(\ell)}_{jn}$, we thus get
$\mathbb{E}\left[S^{(\ell-1)}_{j}\right]=\mathbb{E}\left[N_{\ell-1\to\ell}\right]\mathbb{E}\left[S^{(\ell)}_{j}\right]=\frac{\lambda_{\ell-1}}{\mu_{\ell-1}}\mathbb{E}\left[S^{(j)}_{\ell}\right]=\frac{\lambda_{\ell-1}\cdot\ldots\cdot\lambda_{j-1}}{\mu_{\ell-1}\cdot\ldots\cdot\mu_{j}}=\eta_{\ell-1,j},$
where we have used Wald’s identity in the first step and the induction hypothesis in the penultimate one.
We would like to emphasise that Stefanov (1995) has proved the analogous result for finite-state discrete- or continuous-time birth-death processes, not necessarily absorbing, with the help of analytical properties of certain exponential families. We have complemented this here by a simple probabilistic argument.
Let us also remark that there is a nice intuitive argument for (A.1). If $\lambda_{j}=0$, the transition to $j-1$ at rate $\mu_{j}$ is the only way out of $j$, so the sojourn time in $j$ follows $\mathrm{Exp}(\mu_{j})$ and (A.1) is clear. For $\lambda_{j}>0$, the sojourn time in $j$ may be interrupted by one or several excursions to states $>j$, so the exponential clock for a death event out of state $j$ runs intermittently; but, due to the memoryless property of the exponential distribution, the total sojourn time in $j$ still follows $\mathrm{Exp}(\mu_{j})$, thus (A.1) remains true.
B A derivation of the moments of the path integrals of birth-death processes
The moments (24) of the path integral (20) are usually derived via Laplace transforms (see, for example, Goel and Richter-Dyn (1974, App. D)). Here we take an alternative route that bypasses Laplace transforms and derives a first-step equation in a direct way, thus validating (62) derived in a heuristic way by Norden (1982) via the backward equation.
We first observe, following eq.(83) in Crawford et al. (2018), that our path integral (20) for the birth-death process $(Z(t))_{t\geqslant 0}$ restricted to $[N]_{0}$ with birth and death rates $\lambda_{j}\geq 0$ and $\mu_{j}>0$ for $j\in[N]$, $\lambda_{0}=0$, and the convention $\mu_{0}=\lambda_{N}=0$, equals the first-passge time
$T_{i}^{*}\mathrel{\mathop{:}}=\inf\{t:Z^{*}(t)=0\mid Z^{*}(0)=i\},\quad i\in[N]_{0},$ (B1)
of the “modified" birth-death process $(Z^{*}(t))_{t\geqslant 0}$ with birth and death rates
$\lambda_{j}^{*}\mathrel{\mathop{:}}=\frac{\lambda_{j}}{f(j)},\quad\mu_{j}^{*}\mathrel{\mathop{:}}=\frac{\mu_{j}}{f(j)}\quad\text{for }j\in[N],\quad\text{and}\quad\lambda_{0}^{*}=\mu_{0}^{*}=0$ (B2)
(this is well defined since $f(j)>0$ for $j\in[N]$). Note that, due to the finite state space, $T_{i}^{*}$ is almost surely finite.
Let us now calculate the $m$-th moment of $T_{i}^{*}$, following the standard first-step approach outlined, for example, in (Pinsky and Karlin, 2011, Sec. 6.5.2). Suppose that we are currently in state $i\in[N]$. The sojourn time in this state, whose realisation is denoted by $r$ below, follows the exponential distribution with parameter $\lambda_{i}^{*}+\mu_{i}^{*}$. After the transition away from $i$, the process experiences a waiting time until absorption that is distributed as $T_{i+1}^{*}$ or $T_{i-1}^{*}$, depending on whether it moves from $i$ to $i+1$ (probability $\lambda_{i}^{*}/(\lambda_{i}^{*}+\mu_{i}^{*})$) or $i-1$ (probability $\mu_{i}^{*}/(\lambda_{i}^{*}+\mu_{i}^{*})$). Thus we obtain
$\begin{split}\mathbb{E}\left[(T_{i}^{*})^{m}\right]&=\int_{0}^{\infty}\left\{\frac{\lambda_{i}^{*}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathbb{E}\left[(T_{i+1}^{*}+r)^{m}\right]+\frac{\mu_{i}^{*}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathbb{E}\left[(T_{i-1}^{*}+r)^{m}\right]\right\}(\lambda_{i}^{*}+\mu_{i}^{*})e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\\ &=\lambda_{i}^{*}\mathbb{E}\left[\int_{0}^{\infty}(T_{i+1}^{*}+r)^{m}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\right]+\mu_{i}^{*}\mathbb{E}\left[\int_{0}^{\infty}(T_{i-1}^{*}+r)^{m}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\right]\end{split}$ (B3)
(products containing the factor $\lambda_{N}^{*}=0$ are 0 because $\lambda_{N}=\lambda_{N}^{*}=0$). The integrals in the second line can be evaluated, by integration by parts, as
$\begin{split}\int_{0}^{\infty}(T_{i\pm 1}^{*}+r)^{m}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r&=\left[-(T_{i\pm 1}^{*}+r)^{m}\frac{e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}}{\lambda_{i}^{*}+\mu_{i}^{*}}\right]_{0}^{\infty}+m\int_{0}^{\infty}(T_{i\pm 1}^{*}+r)^{m-1}\frac{e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathrm{d}r\\ &=\frac{(T_{i\pm 1}^{*})^{m}}{\lambda_{i}^{*}+\mu_{i}^{*}}+\frac{m}{\lambda_{i}^{*}+\mu_{i}^{*}}\int_{0}^{\infty}(T_{i\pm 1}^{*}+r)^{m-1}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r,\end{split}$ (B4)
so $\mathbb{E}\left[(T_{i}^{*})^{m}\right]$ turns into
$\begin{split}&\mathbb{E}\left[(T_{i}^{*})^{m}\right]=\frac{\lambda_{i}^{*}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathbb{E}\left[(T_{i+1}^{*})^{m}\right]+\frac{\mu_{i}^{*}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathbb{E}\left[(T_{i-1}^{*})^{m}\right]\\ &\phantom{=}+\frac{m}{\lambda_{i}^{*}+\mu_{i}^{*}}\left\{\lambda_{i}^{*}\mathbb{E}\left[\int_{0}^{\infty}(T_{i+1}^{*}+r)^{m-1}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\right]+\mu_{i}^{*}\mathbb{E}\left[\int_{0}^{\infty}(T_{i-1}^{*}+r)^{m-1}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\right]\right\}.\end{split}$ (B5)
Since the expression in curly brackets equals $\mathbb{E}\left[(T_{i}^{*})^{(m-1)}\right]$ by (B3), we obtain the first-step equation
$(\lambda_{i}^{*}+\mu_{i}^{*})\mathbb{E}\left[(T_{i}^{*})^{m}\right]=\lambda_{i}^{*}\mathbb{E}\left[(T_{i+1}^{*})^{m}\right]+\mu_{i}^{*}\mathbb{E}\left[(T_{i-1}^{*})^{m}\right]+m\mathbb{E}\left[(T_{i}^{*})^{(m-1)}\right],\quad m\geq 1,i\in[N],$ (B6)
with boundary conditions $\mathbb{E}\left[(T_{0}^{*})^{m}\right]=0$ for all $m\geq 1$ and $\mathbb{E}\left[(T_{i}^{*})^{0}\right]=1$ for all $i\in[N]$.
Since (B6) expresses $m$th moments in terms of $(m-1)$st moments, we can now generalise Norden (1982) and recursively solve (B6) from lower to higher moments. To this end, let $\boldsymbol{b}^{(m)}\mathrel{\mathop{:}}=\left(\mathbb{E}\left[(T_{1}^{*})^{m}\right],\cdots,\mathbb{E}\left[(T_{N}^{*})^{m}\right]\right)^{\top}$ for $m\geq 0$; note that $\boldsymbol{b}^{(0)}=\boldsymbol{1}$, a vector whose components are all one. Then (B6) reads
$\boldsymbol{A}^{*}\boldsymbol{b}^{(m)}=m\boldsymbol{b}^{(m-1)},\quad m>0,$ (B7)
where $\boldsymbol{A}^{*}=(A^{*}_{ij})_{i,j\in[N]}$ is an $N\times N$ matrix with elements
${A}^{*}_{ij}=\begin{cases}\lambda_{i}^{*}+\mu_{i}^{*},&j=i,\\ -\lambda_{i}^{*},&j=i+1,\\ -\mu_{i}^{*},&j=i-1,\\ 0,&\text{otherwise}.\end{cases}$ (B8)
It is also clear from (B2) that $\boldsymbol{A}^{*}=\boldsymbol{F}^{-1}\boldsymbol{A}$, where $\boldsymbol{F}$ is the (invertible) $N\times N$ diagonal matrix $\boldsymbol{F}=\textrm{diag}\left[f(1),\cdots,f(N)\right]$ and $\boldsymbol{A}=(A_{ij})_{i,j\in[N]}$ is defined as $\boldsymbol{A}^{*}$ with $\lambda_{i}^{*}$ and $\mu_{i}^{*}$ replaced by $\lambda_{i}^{*}$ and $\mu_{i}^{*}$, respectively. Therefore, (B7) turns into
$\left(\boldsymbol{F}^{-1}\boldsymbol{A}\right)\boldsymbol{b}^{(m)}=m\boldsymbol{b}^{(m-1)}$ (B9)
or, equivalently,
$\boldsymbol{b}^{(m)}=m\left(\boldsymbol{A}^{-1}\boldsymbol{F}\right)\boldsymbol{b}^{(m-1)}=\cdots=m!\left(\boldsymbol{A}^{-1}\boldsymbol{F}\right)^{m}\boldsymbol{b}^{(0)}=m!\left(\boldsymbol{A}^{-1}\boldsymbol{F}\right)^{m}\boldsymbol{1}.$ (B10)
From the standard theory of Markov chains (Kemeny and Snell 1960, Chap. III; see also Norden 1982, p. 694), we know that $\boldsymbol{A}$ is invertible and $\boldsymbol{A}^{-1}$ has elements $\left(\boldsymbol{A}^{-1}\right)_{ij}=\zeta_{ij}$ with $\zeta_{ij}$ of (22); so (24) is an immediate consequence.
C A derivation of the first two moments of the stationary tree length, $L_{n}$
C.1 First moment of $L_{n}$
Inserting (26) into (21) with $f=\mathrm{id}$ leads to
$\mathbb{E}\left[L_{n}\right]=\sum_{j=1}^{N}\sum_{\ell=1}^{\min(n,j)}\frac{1}{u}\left(\frac{s}{Nu}\right)^{j-\ell}(N-\ell)^{\underline{j-\ell}}.$ (C1)
Substituting $m=j-\ell$ and changing summation turns this into
$\mathbb{E}\left[L_{n}\right]=\sum_{m=0}^{N-1}\sum_{\ell=1}^{\min(n,N-m)}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m}(N-\ell)^{\underline{m}}.$ (C2)
The latter expression is further simplified by using
$\sum_{i=a}^{b}i^{\underline{m}}=\frac{(b+1)^{\underline{m+1}}-a^{\underline{m+1}}}{m+1},$ (C3)
which holds for $a,b,m\in\mathbb{N}$ with $a\leq b$ (see Chapter 2 of Graham et al. 2003). In our case, set $a=N-M$ with $M=\min(n,N-m)$ and $b=N-1$. If $n\leq N-m$, then $(N-M)^{\underline{m+1}}=(N-n)^{\underline{m+1}}$. If, on the other hand, $n>N-m$, we have $(N-M)^{\underline{m+1}}=m^{\underline{m+1}}=0$, but also $(N-n)^{\underline{m+1}}=0$ because $m>N-n$. So, in any case, $(N-M)^{\underline{m+1}}=(N-n)^{\underline{m+1}}$, and
$\sum_{\ell=1}^{\min(n,N-m)}(N-\ell)^{\underline{m}}=\sum_{i=N-M}^{N-1}i^{\underline{m}}=\frac{N^{\underline{m+1}}-(N-M)^{\underline{m+1}}}{m+1}=\frac{N^{\underline{m+1}}-(N-n)^{\underline{m+1}}}{m+1},$ (C4)
which allows us to carry out the second sum in (C2) and yields
$\mathbb{E}\left[L_{n}\right]=\sum_{m=0}^{N-1}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m}\frac{N^{\underline{m+1}}-(N-n)^{\underline{m+1}}}{m+1}.$ (C5)
C.2 Second moment of $L_{n}$
A similar calculation enables us to derive the second moment. Using (24) and (26) with $f=\textrm{id}$, we have
$\begin{split}\mathbb{E}\left[L_{n}^{2}\right]&=2\sum_{j_{1}=1}^{N}\sum_{\ell_{1}=1}^{\min(n,j_{1})}\frac{1}{u}\left(\frac{s}{Nu}\right)^{j_{1}-\ell_{1}}(N-\ell_{1})^{\underline{j_{1}-\ell_{1}}}\sum_{j_{2}=1}^{N}\sum_{\ell_{2}=1}^{\min(j_{1},j_{2})}\frac{1}{u}\left(\frac{s}{Nu}\right)^{j_{2}-\ell_{2}}(N-\ell_{2})^{\underline{j_{2}-\ell_{2}}}\\ &=2\sum_{m_{1}=0}^{N-1}\sum_{\ell_{1}=1}^{\min(n,N-m_{1})}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m_{1}}(N-\ell_{1})^{\underline{m_{1}}}\sum_{m_{2}=0}^{N-1}\sum_{\ell_{2}=1}^{\min(m_{1}+\ell_{1},N-m_{2})}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m_{2}}(N-\ell_{2})^{\underline{m_{2}}},\end{split}$ (C6)
where we have set $m_{1}=j_{1}-\ell_{1}$ and $m_{2}=j_{2}-\ell_{2}$ in the second step. By applying (C4) in the first and the last step, it is further evaluated to
$\begin{split}\mathbb{E}\left[L_{n}^{2}\right]&=2\sum_{m_{1}=0}^{N-1}\sum_{\ell_{1}=1}^{\min(n,N-m_{1})}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m_{1}}(N-\ell_{1})^{\underline{m_{1}}}\\ &\qquad\times\sum_{m_{2}=0}^{N-1}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m_{2}}\frac{N^{\underline{m_{2}+1}}-\{N-(m_{1}+\ell_{1})\}^{\underline{m_{2}+1}}}{m_{2}+1}\\ &=2\sum_{m_{1}=0}^{N-1}\sum_{m_{2}=0}^{N-1}\left(\frac{1}{u}\right)^{2}\left(\frac{s}{Nu}\right)^{m_{1}+m_{2}}\\ &\qquad\times\sum_{\ell_{1}=1}^{\min(n,N-m_{1})}\frac{(N-\ell_{1})^{\underline{m_{1}}}N^{\underline{m_{2}+1}}-(N-\ell_{1})^{\underline{m_{1}}}\{N-(m_{1}+\ell_{1})\}^{\underline{m_{2}+1}}}{m_{2}+1}\\ &=2\sum_{m_{1}=0}^{N-1}\sum_{m_{2}=0}^{N-1}\left(\frac{1}{u}\right)^{2}\left(\frac{s}{Nu}\right)^{m_{1}+m_{2}}\sum_{\ell_{1}=1}^{\min(n,N-m_{1})}\frac{(N-\ell_{1})^{\underline{m_{1}}}N^{\underline{m_{2}+1}}-(N-\ell_{1})^{\underline{m_{1}+m_{2}+1}}}{m_{2}+1}\\ &=2\sum_{m_{1}=0}^{N-1}\sum_{m_{2}=0}^{N-1}\left(\frac{1}{u}\right)^{2}\left(\frac{s}{Nu}\right)^{m_{1}+m_{2}}\\ &\qquad\times\left[\frac{\{N^{\underline{m_{1}+1}}-(N-n)^{\underline{m_{1}+1}}\}N^{\underline{m_{2}+1}}}{(m_{1}+1)(m_{2}+1)}-\frac{N^{\underline{m_{1}+m_{2}+2}}-(N-n)^{\underline{m_{1}+m_{2}+2}}}{(m_{1}+m_{2}+2)(m_{2}+1)}\right].\\ \end{split}$ (C7)
D Derivation of the popularity spectrum
Recall that ${P_{l}}$ is the expected number of traits carried by exactly $l$ individuals in an equilibrium population of size $N$. We calculate $(P_{l})_{l\in[N]}$ by adapting the method of Aoki (2018) (see also Strimling et al. (2009); Fogarty et al. (2015, 2017)) to continuous time as follows. Death occurs at rate $u$ to every individual, and when there are $l$ individuals carrying the trait (that is, the trait has popularity $l$), this trait loses a carrier at death rate $d_{l}=ul$, upon which its popularity decreases by one. Every individual that carries a trait with popularity $l$ transmits this trait to some individual that lacks the trait at rate $s(N-l)/N$, so the trait gains a carrier and its popularity increases by one; altogether, therefore, we have birth events at rate $b_{l}=sl(N-l)/N$ when in state $l\in[N]$. Innovation events produce new traits of popularity 1 at rate $N\mu$ without affecting any other trait. Therefore, the $P_{l}$ obey the following system of differential equations:
$\begin{split}\dot{P}_{1}&=N\mu+d_{2}P_{2}-(b_{1}+d_{1})P_{1},\\ \dot{P}_{l}&=b_{l-1}P_{l-1}+d_{l+1}P_{l+1}-(b_{l}+d_{l})P_{l},\quad 2\leq l\leq N,\end{split}$ (D1)
with the convention $d_{N+1}=P_{N+1}=0$; note that the structure is the same as that of the Kolmogorov forward equation for a birth-death process with immigration to state 1. The (unique) stationary solution is given by $b_{l}P_{l}=d_{l+1}P_{l+1}$, $l\in[N-1]$, together with $N\mu=uP_{1}$, which jointly give
[…]
or, more explicitly,
${P_{l}}=\left(\frac{s}{Nu}\right)^{l-1}\frac{\prod_{i=1}^{l-1}i(N-i)}{l!}{P_{1}}=\left(\frac{s}{Nu}\right)^{l-1}\frac{\mu N^{\underline{l}}}{ul},$ (D3)
which is (39).
We can extend this method to obtain the trait frequency spectrum in a sample of size $n<N$. For a trait with popularity $l$, the number of individuals in a sample of size $n$ that carry the trait follows a hypergeometric distribution; more precisely, the probability that $i$ out of the $n$ individuals carry the trait is $\binom{l}{i}\binom{N-l}{n-i}/\binom{N}{n}$. If $C_{n,i}$ is the number of traits carried by exactly $i$ individuals in a sample of size $n$, then the trait frequency spectrum in a sample of size $n$ results as
$\mathbb{E}\left[C_{n,i}\right]=\sum_{l=1}^{N}P_{l}\frac{\binom{l}{i}\binom{N-l}{n-i}}{\binom{N}{n}},$ (D4)
which is (40). As a consistency check, note that
$\mathbb{E}\left[C_{n}\right]=\sum_{i=1}^{n}\mathbb{E}\left[C_{n,i}\right]$ (D5)
by definition, and, since
$\sum_{i=1}^{n}\frac{\binom{l}{i}\binom{N-l}{n-i}}{\binom{N}{n}}=1-\frac{\binom{l}{0}\binom{N-l}{n}}{\binom{N}{n}}=1-\frac{(N-l)^{\underline{n}}}{N^{\underline{n}}}=1-\frac{(N-n)^{\underline{l}}}{N^{\underline{l}}}$
(where the first step comes from the normalisation of the hypergeometric distribution and the last step is true because $(N-l)^{\underline{n}}N^{\underline{l}}=N^{\underline{n+l}}=(N-n)^{\underline{l}}N^{\underline{n}}$)), (D4) and (D5) together give
$\mathbb{E}\left[C_{n}\right]=\sum_{l=1}^{N}P_{l}\left(1-\frac{(N-n)^{\underline{l}}}{N^{\underline{l}}}\right)=\mu\mathbb{E}\left[L_{n}\right],$ (D6)
where the last step uses (D3) and (27); the result agrees with $\mathbb{E}\left[C_{n}\right]$ of (18). Specifically, for $n=N$, we get $\mathbb{E}\left[C_{N}\right]=\sum_{l=1}^{N}P_{l}$ for the expected total number of distinct traits in the population (denoted by $C_{\rm{pop}}$ in some of the previous studies), as it must.